### v1

# From Equations to Insights: Unraveling Symbolic Structures in PDEs with LLMs

Abstract
Motivated by the remarkable success of artificial intelligence (AI) across diverse fields, the application of AI to solve scientific problems—often formulated as partial differential equations (PDEs)—has garnered increasing attention. While most existing research concentrates on theoretical properties (such as well-posedness, regularity, and continuity) of the solutions, alongside direct AI-driven methods for solving PDEs, the challenge of uncovering symbolic relationships within these equations remains largely unexplored. In this paper, we propose leveraging large language models (LLMs) to learn such symbolic relationships. Our results demonstrate that LLMs can effectively predict the operators involved in PDE solutions by utilizing the symbolic information in the PDEs. Furthermore, we show that discovering these symbolic relationships can substantially improve both the efficiency and accuracy of the finite expression method for finding analytical approximation of PDE solutions, delivering a fully interpretable solution pipeline. This work opens new avenues for understanding the symbolic structure of scientific problems and advancing their solution processes.
Keywords Large Language Models  $\cdot$ Finite Expression Method  $\cdot$ High Dimensional PDEs  $\cdot$ Operator Relation
1 Introduction
Numerous grand challenges in scientific fields, including physics, engineering, biology, and chemistry, can be modeled as partial differential equations (PDEs), which relate unknown functions defined on a multidimensional domain (e.g., in terms of space and time) to its rates of change with respect to each of those variables (Evans, 2022). Mathematically, a PDE takes the following form
$Du(x)=f(u(x),x),\quad x\in\Omega,\quad Bu(x)=g(x),\quad x\in\partial\Omega,$ (1)
where $D$ denotes a certain differential operator defining the dynamics, $Bu=g$ represents the boundary condition (such as Dirichlet, Neumann, and Robin), and $\Omega\subseteq\mathbb{R}^{d}$ describes the domain with boundary $\partial\Omega$.
PDEs often arise in high-dimensional settings (i.e., $d$ is large); for example, well-known instances such as the Poisson equation (Yu et al., 2018), the linear conservation law (Chen et al., 2021), and the nonlinear Schrödinger equation (Han et al., 2020) regularly involve multiple spatial and temporal variables. While low-dimensional PDEs can be solved reliably and efficiently via traditional mesh-dependent methods such as finite difference methods (LeVeque, 2007), finite elements methods (Hughes, 2003), finite volume methods (Toro, 2013), and spectral methods (Boyd, 2001), developing efficient and accurate algorithmic frameworks for computing numerical solutions to high-dimensional PDEs remains an important and challenging topic due to the curse of dimensionality (i.e., the computational costs grow exponentially with respect to the dimensionality of the problem’s domain) (E et al., 2021).
Before solving a PDE, one must understand certain fundamental properties of the solution $u(x)$ that follow from the PDE itself and from the data—namely, $f(u(x),x)$ and $g(x)$. These considerations align with the classical theory of PDEs (Evans, 2022), which focuses on establishing well-posedness (i.e., the existence and uniqueness of solutions) and their regularity and continuity. However, the underlying symbolic relationship among $u$, $f$ and $g$ often remains underexplored. To address this gap, we pose the following question:
Assuming that the PDE (1) admits an analytical solution $u$, and given the operators appearing in the expressions of $f$ and $g$, what are the corresponding operators in the expression for $u$?
Understanding this symbolic connection is crucial for deriving fully interpretable, closed-form solutions, a feature lacking in most existing solution methods, particularly in the high-dimensional setting. As an illustrative example, we consider solving the PDE (1) via the least square method (Dissanayake and Phan-Thien, 1994; Lagaris et al., 1998; Sirignano and Spiliopoulos, 2018), which defines a straightforward functional to characterize the error of the estimated solution by
$L(u)=\|Du-f\|^{2}_{L^{2}(\Omega)}+\lambda\|Bu-g\|^{2}_{L^{2}(\partial\Omega)},$ (2)
where $\lambda>0$ balances the influence of boundary conditions. The goal is to find a function $u^{*}\in L^{2}(\Omega)$ so that the least square loss is minimized, i.e.,
$u^{*}=\mathrm{argmin}_{u\in L^{2}(\Omega)}\;L(u).$ (3)
Because the search space $L^{2}(\Omega)$ is an infinite dimensional space, finding the solution $u^{*}$ is extremely challenging. Fortunately, the search space can be significantly simplified if one knows the operators appearing in the solution in advance, making the search of the optimal solution much easier and effective. This can be accomplished by adopting the recently developed finite expression method (FEX) (Liang and Yang, 2022).
To uncover the symbolic relationship described above, we leverage large language models (LLMs) as predictive tools. The process begins with the generation of a comprehensive, structured dataset composed of symbolic expressions derived from a diverse range of PDE types, including elliptic, parabolic, and hyperbolic equations, along with their associated boundary and initial conditions. These symbolic expressions are represented as computational trees (see Figure 1), where nodes are populated with unary (e.g., sin, cos, exp) and binary operators (e.g., +, -, *, /). This tree structure captures the hierarchical nature of mathematical expressions and ensures a systematic representation of the PDE solutions.
To prepare the data for LLM processing, the symbolic expressions are tokenized and converted into postfix notation (Reverse Polish Notation). This format is chosen because it aligns well with the sequential input requirements of LLMs and eliminates the need for parentheses, simplifying the parsing process. More importantly, it reflects the relative positional relationship of operators in the expression. The tokenized dataset is then used to fine-tune foundational LLMs such as T5, BART, and Llama. During fine-tuning, the models are trained to predict the sequence of operators that constitute the symbolic solution to a new PDE. The training objective is to maximize the accuracy of operator sequence prediction, ensuring that the model can reliably identify the minimal set of operators required to represent the true solution.
Once fine-tuned, the LLM acts as a predictive tool to generate operator sets for new PDE problems. The predicted operator sets are highly refined, containing only the operators necessary to construct the solution, thereby eliminating redundant or irrelevant operators. This refinement is crucial for the subsequent application of the finite-expression method (FEX), a reinforcement learning based symbolic regression technique that searches for the optimal combination of operators and constants to approximate the PDE solution. By providing FEX with a reduced and targeted operator set, the combinatorial search space is significantly narrowed. This reduction not only enhances the accuracy of the solution by focusing on relevant operators, but also dramatically improves computational efficiency by reducing the number of potential combinations that need to be evaluated.
In summary, the integration of LLMs into this workflow transforms the process of symbolic PDE solving. By generating structured datasets, fine-tuning models to predict operator sequences, and leveraging these predictions to guide FEX, we achieve a more accurate and computationally efficient approach to discovering symbolic solutions for complex PDEs. This methodology bridges the gap between machine learning and symbolic mathematics, offering great potential for scientific computing and engineering applications. Particularly, our contributions are summarized as follows:
(1) We formally pose the problem of discovering the symbolic relationship between the solution of a PDE and the associated problem data, including the boundary conditions and right-hand-side functions. Establishing this connection can provide deeper insights into the underlying structure of PDEs, enabling more precise analytical solutions and facilitating the development of efficient numerical methods which provides high-quality analytical solutions. By highlighting the potential impact of such relationships, our work lays the foundation for further studies in this area, aiming to bridge the gap between problem formulation and solution representation.
(2) We propose fine-tuning LLMs to learn the symbolic relationship between problem data and the solution of a PDE, demonstrating remarkable effectiveness in predicting the operators that characterize the solution. By integrating this approach with the FEX method, our numerical findings show that while LLMs may not be inherently suited for directly solving PDEs or other complex scientific problems, they provide substantial benefit in identifying underlying structures and symbolic patterns. This capability significantly streamlines and enhances the solution process. Our work underscores the potential of LLMs as transformative tools in the analytical and computational study of PDEs, bridging the gap between symbolic reasoning and sophisticated mathematical problem-solving.
1.1 Related work
Artificial intelligence (AI) for Science takes advantage of modern advanced machine learning (ML) methods to facilitate, accelerate and enhance scientific discovery (Stevens et al., 2020). Its importance lies in the ability to handle massive datasets, uncover hidden patterns, and simulate complex phenomena. Inspired by the remarkable success of AI across various domains, deep neural network (DNN)-based approaches (LeCun et al., 2015; Goodfellow, 2016) have become increasingly popular for solving (high-dimensional) PDEs. These methods offer certain appealing advantages over traditional numerical techniques, making them a compelling alternative for tackling complex problems. First, DNN-based approaches empirically mitigate the curse of dimensionality that plagues conventional discretization approaches, leading to more efficiency in handling high-dimensional PDEs Han et al. (2018). Second, DNNs, acting as mesh-free universal function approximators (Cybenko, 1989; Hornik et al., 1989; Goodfellow, 2016), can be trained to learn highly complex PDE solutions without the need of prespecifying basis functions or meshes (Berg and Nyström, 2018; Raissi et al., 2019). Hence, they are highly suitable to approximate complex solution spaces that challenges the traditional mesh-based approaches. Third, the solutions represented by DNNs can be evaluated efficiently once trained, which turns out to be a crucial property for tasks, including uncertainty quantification and real-time optimal control, require calling PDE solutions repeatedly (Zhu et al., 2019). Last but not least, continuous progress in designing comprehensive neural network architectures—such as Fourier neural operators and advanced physics-informed neural network frameworks—alongside theoretical advancements in understanding their approximation capabilities, are making these methods increasingly applicable (Lu et al., 2019; Li et al., 2020).
However, several critical factors often hinder deep neural networks (DNNs) from achieving highly accurate solutions, even for relatively simple problems. These include the need for large, diverse, and high-quality training data (Szegedy, 2013; Buda et al., 2018), sensitivity to hyperparameter selection (Bergstra and Bengio, 2012), the challenge of optimizing highly nonconvex objectives (Dauphin et al., 2014; Choromanska et al., 2015), issues like vanishing and exploding gradients (Glorot and Bengio, 2010), and poor generalization capabilities (Zhang et al., 2021). Additionally, solutions produced by DNNs often lack interpretability, preventing users from leveraging insights for future decision-making (Rudin, 2019). To overcome these challenges and achieve both highly accurate and interpretable solutions for PDEs, Liang and Yang (2022) recently introduced the finite expression method (FEX). This approach leverages recent advances in approximation theory, which demonstrate that high-dimensional functions can be effectively approximated using functions composed of simple operators (Shen et al., 2021a, b; Zhang et al., 2022). This methodology seeks approximate PDE solutions within the space of functions composed of finitely many analytic expressions, effectively avoiding the curse of dimensionality. In FEX, the mathematical expression representing the PDE solution is modeled as a binary tree, where each node holds either a unary or binary operator chosen from pre-specfied operator sets. The goal is to find the optimal sequence of operators and coupling parameters that minimize the objective function associated with the PDE. This leads to a combinatorial optimization (CO) problem, which is then solved by a deep reinforcement learning (RL) method with the policy gradient optimizer. The FEX outperforms certain existing numerical PDE solvers in terms of accuracy, interpretability and memory efficiency, making it a promising approach for solving high-dimensional PDEs and other complex problems; see e.g., (Song et al., 2023, 2024; Hardwick et al., 2024).
2 Methodology
In this section, we propose a novel method that fine-tunes large language models (LLMs) to predict the operator sets present in the symbolic expressions of partial differential equation (PDE) solutions. Our approach consists of three main stages: data generation, model fine-tuning, and performance evaluation. Below, we detail the pipeline for generating synthetic data, the procedure for fine-tuning the LLMs, and the evaluation of their predictive capabilities.
2.1 Binary Computational Trees for PDE Expressions
Expressions in PDEs are represented as binary computational trees, as depicted in Figure 1. These trees provide a structured and hierarchical way to encode mathematical expressions, enabling efficient exploration and manipulation of the solution space. The tree generation process is systematic and begins by specifying the depth of the binary tree, which determines the complexity and granularity of the resulting expressions.
Figure 1: Computational expression tree (Liang and Yang, 2022).
The construction of the tree involves the following steps:
  1. 1.
    Depth Specification: The depth of the tree is predefined, controlling the number of hierarchical levels in the tree. A deeper tree allows for more complex expressions but increases computational cost.
  2. 2.
    Recursive Construction: At each level of depth, the tree is built recursively. Starting from the root node, each node is expanded by adding child nodes according to the rules defined for operators. This recursive process ensures a systematic exploration of the solution space, covering a wide range of possible expressions.
Each node in the computational tree represents an operator, which can be either a unary operator or a binary operator:
  • Unary Operators: These operators act on a single operand and include functions such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other nonlinear transformations ($\sqrt{\cdot}$, ||$\cdot$|| ⋅ |, $(\cdot)^{k}$, $k\in\mathbb{Z}$). For each node with a unary operator, two scalars ($\alpha$ and $\beta$) are coupled with the node. These scalars enable the composition of the underlying operation with an affine function, resulting in expressions of the form $\alpha\cdot u(\cdot)+\beta$, where $u(\cdot)$ is the unary operation. This coupling ensures that the tree can represent both nonlinear transformations and linear adjustments, enhancing its expressive power.
  • Binary Operators: These operators require two operands and include standard arithmetic operations including addition ($+$), subtraction (--), multiplication ($\times$) and division (///). Binary operators combine two sub-expressions (left and right child nodes) to form more complex expressions. For example, a multiplication node might combine the results of two subtrees representing different terms in an expression.
The use of binary computational trees offers several advantages: (1) The tree structure provides a clear and interpretable representation of mathematical expressions, making it easier to analyze and manipulate them; (2) By recursively constructing trees of varying depths, the method ensures a thorough exploration of the solution space, capturing both simple and complex expressions; and (3) The inclusion of both unary and binary operators, along with affine transformations, allows the tree to represent a wide range of mathematical operations and relationships. By systematically generating and evaluating computational trees, it becomes possible to identify candidate expressions that best describe the underlying dynamics of a system. Additionally, the framework can be combined with advanced techniques, including optimization and reinforcement learning, to refine the scalars and operators, further improving the accuracy of the discovered models.
2.2 Data Generation for Equation Types
In this subsection, we describe the data generation process for constructing a structured dataset of symbolic expressions derived from various PDE types. The dataset is designed to capture the symbolic relationships between the solution of a PDE and the associated problem data, including the boundary conditions and right-hand-side functions. The data generation process is outlined in Figure 2.
Figure 2: Data generation pipeline.
The construction of the one data point involves the following steps.
PDE Type Specification. We first specify the type of PDE (e.g., Poisson, linear conservation law, and Schrödinger) and the type of boundary conditions (e.g., Dirichlet, Neumann and Robin) based on the information of the operators $D$ and $B$. This results in the first component of the data point, which includes the PDE type and boundary condition type of the data point.
Randomly generated function $u(x)$. Next, we generate a random function $u(x)$ by constructing a binary computational tree with a specified depth, as described in Section 2.1. The tree is populated with random unary and binary operators drawn from prescribed operator sets that are rich enough. Consequently, we obtain the expression tree for the solution function $u(x)$, which is used to derive the second component of the data point, i.e., expressions of the right-hand-side function $f$ and the function $g$ defined on the boundary of the domain.
Postfix Representation. After obtaining the expressions of the functions $u(x)$, $f(u(x),x)$, and $g(x)$, we convert them into postfix notation. This conversion simplifies the tokenization process and aligns the expressions with the input format required by LLMs. The postfix notation also eliminates the need for parentheses, making the expressions more interpretable and easier to parse. More importantly, it reflects the relative positional relationship of operators in the expression, which is crucial for the LLMs to learn the symbolic relationships between the operators. In the end, we obtain the second component of the data point, which includes the operators used in the right-hand-side function $f(u(x),x)$, and the function $g(x)$. These input component are then concatenated to generate the input. Meanwhile, the label of the data set is obtained similarly by converting the expression of the function $u(x)$ into postfix notation. See Figure 3 for the detailed convertion pipeline and a simple illustrated example.
Figure 3: Pipeline for converting data into postfix notation and an illustrated example.
2.3 Fine-tuning
Inspired by the success of LLMs, we propose fine-tuning the sequence-to-sequence Transformer architecture that leverages cross-attention to incorporate PDE context when generating solution operators. This architecture is selected because it effectively captures long-range dependencies and the complex hierarchical structure inherent in PDE metadata, thereby learning an effective mapping between symbolic operator sequences.
We first preprocess and tokenize the dataset generated as in the previous subsection, and then fine-tune the LLM using the proceed data using the common practices in the literature. The whole pipeline is outlined in Figure 4 and the key components in the pipeline are discussed in detailed as follows.
Figure 4: Pipeline for tokenization and embedding of operator sequences for LLM fine-tuning
Tokenization. We begin by loading the PDE data in jsonl format, shuffling it, and splitting it into training and validation sets. Each data point is then processed by a textual embedding layer that converts operator symbols (e.g., sin, cos, exp) into tokens. To ensure a robust and consistent representation, we use HuggingFace’s  [^1] https://github.com/huggingface AutoTokenizer, which selects an appropriate tokenization scheme based on the underlying pre-trained model. The tokenizer pads sequences to a fixed length and incorporates learnable positional embeddings. Additionally, an attention mask is generated to differentiate between valid and padding tokens, facilitating self-attention computations. Finally, padding tokens in the label sequences are replaced with $-100$, the default ignore_index in PyTorch’s  [^2] https://pytorch.org/ nn.CrossEntropyLoss, ensuring they are ignored during loss computation. This approach standardizes the transformation of textual operator representations into numerical tokens for model training.
Architecture and Finetuning: Next, we select a pretrained LLM with a sequence-to-sequence Transformer architecture. We fine-tune this model using the HuggingFace Trainer API, which provides a suite of tools for efficient optimization. During fine-tuning, we minimize the token-wise cross-entropy loss, which penalizes incorrect token predictions, including misclassification of the End-of-Sequence (EOS) token. Backpropagation is performed to update the model’s parameters according to the optimization algorithm. We emphasize the use of teacher forcing  [^3] Teacher forcing is a training strategy where, at each time step, the decoder is provided with the actual ground-truth token instead of its own previous prediction. during fine-tuning. By supplying the decoder with the full target sequence, this approach ensures that the token-wise cross-entropy loss is computed over aligned positions, preventing error accumulation and reinforcing the correct sequential dependencies more effectively and robustly.
2.4 Inferencing
Once trained, the model can predict operator sets in solutions for new PDEs. Specifically, we provide the language model with a data point consisting of the PDE type, its right-hand side (RHS), and the associated boundary condition, as described in Section 2.2. This data is then tokenized following the procedure outlined in the previous subsection. The language model’s decoder then auto-regressively generates an operator sequence representing the solution, with each predicted token conditioned on previously decoded tokens and the input context. While the exact sequential order may vary slightly, the model consistently identifies the correct set of operators, prioritizing presence over strict ordering.
Since the model’s output is a numerical vector, a final post-processing step is required to extract the predicted operators. First, we extract the unique operators by removing duplicates from the predicted postfix representation, ensuring that only essential components remain. Because the model is trained to reproduce entire operator sequences rather than direct sets, this step distills the output into a concise and meaningful form. Next, the obtained set is checked for misspellings and tokenization errors. Given that the model generates sequences based on learned patterns rather than strict syntax rules, occasional character-level errors may arise, particularly with uncommon symbols or operators. To ensure correctness and consistency, any malformed operators are discarded.
To evaluate the accuracy of the fine-tuned LLM, we use the squared norm to quantify the discrepancy between the predicted operator set and the ground-truth label. Particularly, given a dictionary of all possible operators of size $n$, each operator set is represented as a binary vector, as illustrated in Table 1.
Binary Vector
Function Preprocessed Operator Set
$x_{1}$
$x_{2}$
^2
^3
SIN
COS
EXP
$(5x_{1})^{2}+\sin(3x_{2})\cdot x_{2}$ [ $x_{1}$, $x_{2}$, ^2, +, *, SIN ] 1 1 1 0 1 1 1 0 0
$5\exp(2x_{1})\cdot(\cos(6x_{1}))^{3}$ [ $x_{1}$, ^3, *, COS, EXP ] 1 0 0 1 0 1 0 1 1
Table 1: Sample binary-vector encoding for two example expressions. The size of the dictionary is 9.
Then, the squared distance between two such operator sets, represented by vectors $y\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{n}$, can be defined as
$\|y-z\|^{2}=\sum_{i=1}^{n}\left(y_{i}-z_{i}\right)^{2}.$
Clearly, it measures the number of mismatched operators between the two operator sets. Moreover, this metric prioritizes the presence of operators over their sequence order, ensuring that the predicted operators remain useful for numerical and analytical PDE solving, as well as for scientific analysis of PDE models.
3 Experimental Results
In this section, we conduct extensive experiments to evaluate the effectiveness of our proposed approach. Our evaluation consists of two main components: (1) assessing the predictive performance of fine-tuned large language models (LLMs) in identifying operator sets within PDE solutions and (2) demonstrating the practical impact of these predictions in guiding operator selection within the finite expression method (FEX).
First, we analyze the ability of fine-tuned LLMs to predict the correct operator sets in PDE solutions. The objective is to assess whether the models can effectively learn and generalize the symbolic relationships between PDE solutions and problem data. We compare the predicted operator sets against ground-truth labels to measure the accuracy of the learned representations. Additionally, we examine the impact of model architecture by fine-tuning several commonly used open-source LLMs.
Second, we apply the fine-tuned LLMs to enhance the efficiency and accuracy of the FEX method by leveraging their predicted operator sets. This experiment attempts to achieve two goals. On one hand, we demonstrate that by obtaining more precise operator sets of smaller sizes, the computational costs spent in FEX can be significantly reduced. On the other hand, we evaluate whether FEX achieves improved accuracy in learning PDE solutions when provided with the correct prior information about the relevant operators. By comparing FEX solutions obtained with and without LLM-guided operator selection, we quantify improvements in both computational efficiency and numerical accuracy.
Through these experiments, we aim to establish the viability of fine-tuned LLMs as effective tools for discovering symbolic relations in PDE solutions and enhancing PDE-solving techniques.
3.1 Experimental Setup
For data generation, we consider two types of PDEs—namely, the Poisson equation and the linear conservation law—each paired with three commonly used boundary conditions: Cauchy, Dirichlet, and Neumann. Using a tree depth of 3, we randomly generate a dataset of 60,000 equations, with 30,000 examples for each PDE type, as described in Section 2.2. This dataset is then used to fine-tune two state-of-the-art large language models: BART (Lewis et al., 2019) and T5 (Raffel et al., 2020).
For fine-tuning, we optimize the cross-entropy loss using the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and a weight decay of $0.01$. The models are trained for 3 epochs with a batch size of 2. Fine-tuning is conducted on an NVIDIA A6000 GPU with 48 GB of memory. To efficiently manage storage, we save checkpoints every 500 steps and retain only the three most recent checkpoints.
Finally, we integrate the model into the FEX solver [^4] For a detailed description of the FEX method, we refer the reader to Liang and Yang (2022), and for the implementation of the original FEX, see https://github.com/LeungSamWai/Finite-expression-method.. During the execution of FEX for a given PDE, the model is invoked for inference. After inference, the post-processing step detailed in Section 2.4 is applied to extract the unique unary and binary operator sets. These sets are then used to configure the available operators within FEX during its search for an analytical PDE solution. We refer to this version of FEX as the "LLM-informed FEX" and denote the original FEX as "uninformed FEX." The uninformed FEX employs a fixed operator set:
[…]
In contrast, the LLM-informed FEX dynamically predicts the operator sets for each example, enabling a more adaptive and potentially efficient search process.
3.2 Effectiveness of Fine-tuned LLMs
In this subsection, we evaluate the performance of the fine-tuned BART and T5 models in predicting operator sets for the two PDE types. Specifically, we track the training and validation losses throughout the training process, as well as the average number of mismatched operators on the test dataset for each epoch. The computational results are summarized in Tables 2 and 3.
Epoch Training Loss Validation Loss Average Mismatch
1 0.058200 0.057948 2.475833
2 0.040000 0.046193 2.020000
3 0.039200 0.038100 1.720210
Table 2: Fine-tuning BART on large dataset.
Epoch Training Loss Validation Loss Average Mismatch
1 0.046800 0.045303 1.711056
2 0.037200 0.036140 1.544056
3 0.012300 0.020392 1.012000
Table 3: Fine-tuning T5 on large dataset.
Both fine-tuned models exhibit a consistent decrease in training and validation loss across epochs, indicating successful fine-tuning. However, T5 demonstrates a more pronounced reduction in both loss metrics. The larger gap between BART’s training and validation loss suggests that it may be more prone to overfitting compared to T5, which appears to generalize better. In terms of the average number of mismatched operators, T5 consistently outperforms BART in this regard across all epochs. The sharper decline in mismatch for T5 suggests that it adapts more effectively to the symbolic representation of PDE solutions and captures the underlying operator structures with higher precision. Overall, T5 demonstrates superior performance in predicting operators for PDE solutions, making it a more suitable choice for integration into the LLM-informed FEX solver. In light of these findings, we incorporate the fine-tuned T5 model into the FEX framework to enhance its operator selection process.
3.3 LLM-informed FEX
In this subsection, we evaluate the effectiveness of the LLM-informed FEX using the fine-tuned T5 model in comparison to the uninformed FEX. As an initial validation step, we conduct a sanity check using several simple examples to ensure the correctness of the approach. To comprehensively assess the performance of the proposed framework, we further test both versions of FEX on a dataset of 100 randomly generated PDEs, consisting of 50 Poisson equations and 50 linear conservation law equations.
3.3.1 Sanity Check
We first construct 4 Poisson equations and 4 linear conservation law equations, whose true solutions, denoted as $u(x)$, are randomly handcrafted. For each example, we present the Binary and Unary set sizes, the number of iterations and computational time taken to find the approximate solution, and the error with respect to the true solution. The results are summarized in Tables 4 and 5, respectively.
$u(x)$ Method Binary Size Unary Size Iters Time [m] Error
$4\,\cos\bigl{(}4x^{2}\cos(x_{0})\bigr{)}$ LLM-informed 1 2 4.25 8.25 0
Uninformed 3 9 167 340 0
$4\,x_{1}^{3}+4\,x_{1}^{2}+2\,\cos\bigl{(}4\,x_{1}^{3}\cos(x_{0})\bigr{)}$ LLM-informed 2 4 102 286 $10^{-8}$
Uninformed [^5] Uninformed FEX did not converge within 2000 iterations, indicated by “2000+” (iterations) and “N/A” for error. 3 9 2000+ 2400+ N/A
$128\,x_{2}^{3}+2\,x_{2}\,e^{4\,x_{2}^{4}}$ LLM-informed 2 4 34.5 40.5 $4\times 10^{-7}$
Uninformed 3 9 90 186 $6\times 10^{-7}$
$64\,x_{1}^{2}\,e^{2\,x_{2}}$ LLM-informed 1 3 15.5 21 $3\times 10^{-8}$
Uninformed 3 9 103.5 161 $3\times 10^{-8}$
Table 4: Sanity check for the Poisson equation.
$u(x)$ Method Binary Size Unary Size Iters Time [m] Error
$\sin\Bigl{(}x_{1}+\tfrac{\pi}{4}x_{2}+\tfrac{\pi}{4}x_{3}\Bigr{)}$ LLM-informed 2 2 1 0.3 0
Uninformed 3 9 3 1 0
$4\,\bigl{(}2x_{0}-2x_{2}^{3}\bigr{)}^{3}$ LLM-informed 2 2 2 0.6 0
Uninformed 3 9 33.3 9.4 0
$8\,x_{2}^{2}+8\,\sin(x_{1})$ LLM-informed 2 3 1 1 0
Uninformed 3 9 24 28.5 0
[…]
Uninformed 3 9 273 488 0
Table 5: Sanity check for the linear conservation law euqation.
The computational results presented in Tables 4 and 5 demonstrate that the LLM-informed FEX significantly improves efficiency over the uninformed FEX while maintaining high accuracy. One major advantage is the reduction in the number of required operators. The uninformed FEX consistently uses a fixed operator set of 3 binary and 9 unary operators, whereas the LLM-informed FEX dynamically selects a smaller, more relevant subset—typically with at most 2 binary and 4 unary operators. This reduced search space leads to a dramatic decrease in the number of iterations required for convergence. For instance, in the first Poisson equation example, the LLM-informed FEX requires only 4.25 iterations compared to 167 for the uninformed version, and in another case, the uninformed FEX exceeds 2000 iterations without converging, while the LLM-informed approach successfully completes the solution in 102 iterations. This efficiency gain translates into significant computational time savings, with the LLM-informed approach solving problems up to an order of magnitude faster. Despite this drastic reduction in computational cost, the accuracy remains comparable; for cases where the uninformed FEX converges, both methods yield similar approximation errors, and for the linear conservation law equations, both achieve exact solutions. These results highlight the effectiveness of integrating the fine-tuned T5 model into the FEX framework, demonstrating that LLM-informed operator selection leads to a more efficient and reliable symbolic PDE solver.
3.3.2 Randomly Generated PDEs
To further validate the reliability of our approach, we evaluated the performance of the FEX algorithm on a set of randomly generated PDE examples.
For each PDE type, we randomly generated 50 instances using a depth-2 computational tree structure, as described in Section 2.2. Since FEX is a randomized reinforcement learning approach, each instance was solved five times to obtain average computational results, including the number of iterations until convergence, computation time, and solution error. The FEX algorithm was limited to a maximum of 150 iterations and an error tolerance of $10^{-3}$ (i.e., $\varepsilon<1\times 10^{-3}$), where $\varepsilon$ was computed during training using a composite loss that combines PDE residuals and boundary violations over a set of sampled points from the domain and its boundary, respectively. A PDE instance was considered successfully solved if the desired accuracy was achieved within 150 iterations. The computational results are summarized in Tables 6 and 7.
Table 6: Summary Statistics for Poisson PDEs
Metric Uninformed FEX LLM-Informed FEX Improvement (Reduction)
Success Ratio (Valid Solutions) 51.06% (24/47) 68.09% (32/47) $\uparrow 33.35\%$
For Cases Where BOTH Versions Yield Valid Solutions
Mean Iterations 61.80 27.87 $\downarrow 54.90\%$
Mean Duration (min) 40.73 20.25 $\downarrow 50.27\%$
Table 7: Summary Statistics for Linear Conservation Law PDEs
Metric Uninformed FEX LLM-Informed FEX Improvement (Reduction)
Success Ratio (Valid Solutions) 66.67% (30/45) 64.44% (29/45) $\downarrow 3.46\%$
For Cases Where BOTH Versions Yield Valid Solutions
Mean Iterations 24.87 13.98 $\downarrow 43.80\%$
Mean Duration (min) 6.43 3.60 $\downarrow 44.03\%$
The results in Tables 6 and 7 indicate that the LLM-informed FEX method consistently accelerates convergence compared to the Uninformed FEX, though its impact on solution validity varies across PDE types. For Poisson PDEs, the LLM-informed approach significantly improves the success ratio, increasing it by 33.35% (from 51.06% to 68.09%), while also reducing the number of iterations by 54.90% and computation time by 50.27%. This suggests that incorporating LLM-informed heuristics enhances both accuracy and efficiency for this class of problems. However, for linear conservation law PDEs, the LLM-informed approach slightly reduces the success ratio by 3.46%, indicating that its effectiveness in improving solution validity may be problem-dependent. This decline appears to stem from the inherent instability of FEX, even when accurate operators are provided. Despite this, the LLM-informed method still achieves a 43.80% reduction in iterations and a 44.03% decrease in computation time, demonstrating substantial gains in computational efficiency.
Overall, the LLM-informed approach consistently accelerates convergence, making it a valuable tool for solving PDEs more efficiently. However, its impact on solution validity remains inconsistent. This suggests that while the LLM-informed approach effectively reduce computational costs, one may need to conduct further refinement to the PDE solver itself to improve robustness across different PDE types.
4 Conclusions
In this work, we explored the potential of large language models (LLMs) to uncover symbolic relationships in partial differential equations (PDEs), a largely unexplored challenge in the intersection of AI and scientific computing. Our results demonstrate that fine-tuned LLMs can effectively predict the operators involved in PDE solutions by leveraging symbolic information from the governing equations. By integrating these predictions into the finite expression method (FEX), we significantly enhanced both the efficiency and accuracy of analytical PDE approximations. Compared to the traditional uninformed FEX, the LLM-informed approach reduces the number of operators required, accelerates convergence, and maintains high solution accuracy, providing a fully interpretable and computationally efficient pipeline. These findings highlight the promising role of LLMs in advancing symbolic reasoning for scientific problems, paving the way for further exploration of AI-driven methodologies in mathematical modeling and equation solving.
Acknowledgement
The authors were partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and the DARPA D24AP00325-00.

### v2

@@
 Abstract
-Motivated by the remarkable success of artificial intelligence (AI) across diverse fields, the application of AI to solve scientific problems—often formulated as partial differential equations (PDEs)—has garnered increasing attention.
+Motivated by the remarkable success of artificial intelligence (AI) across diverse fields, the application of AI to solve scientific problems, often formulated as partial differential equations (PDEs), has garnered increasing attention.
 While most existing research concentrates on theoretical properties (such as well-posedness, regularity, and continuity) of the solutions, alongside direct AI-driven methods for solving PDEs, the challenge of uncovering symbolic relationships within these equations remains largely unexplored.
 In this paper, we propose leveraging large language models (LLMs) to learn such symbolic relationships.
-Our results demonstrate that LLMs can effectively predict the operators involved in PDE solutions by utilizing the symbolic information in the PDEs.
-Furthermore, we show that discovering these symbolic relationships can substantially improve both the efficiency and accuracy of the finite expression method for finding analytical approximation of PDE solutions, delivering a fully interpretable solution pipeline.
+Our results demonstrate that LLMs can effectively predict the operators involved in PDE solutions by utilizing the symbolic information in the PDEs both theoretically and numerically.
+Furthermore, we show that discovering these symbolic relationships can substantially improve both the efficiency and accuracy of symbolic machine learning for finding analytical approximation of PDE solutions, delivering a fully interpretable solution pipeline.
 This work opens new avenues for understanding the symbolic structure of scientific problems and advancing their solution processes.
-Keywords Large Language Models  $\cdot$ Finite Expression Method  $\cdot$ High Dimensional PDEs  $\cdot$ Operator Relation
 1 Introduction
-Numerous grand challenges in scientific fields, including physics, engineering, biology, and chemistry, can be modeled as partial differential equations (PDEs), which relate unknown functions defined on a multidimensional domain (e.g., in terms of space and time) to its rates of change with respect to each of those variables (Evans, 2022).
+Numerous grand challenges in scientific fields, including physics, engineering, biology, and chemistry, can be modeled as partial differential equations (PDEs), which relate unknown functions defined on a multidimensional domain (e.g., in terms of space and time) to its rates of change with respect to each of those variables [11].
 Mathematically, a PDE takes the following form
-$Du(x)=f(u(x),x),\quad x\in\Omega,\quad Bu(x)=g(x),\quad x\in\partial\Omega,$ (1)
-where $D$ denotes a certain differential operator defining the dynamics, $Bu=g$ represents the boundary condition (such as Dirichlet, Neumann, and Robin), and $\Omega\subseteq\mathbb{R}^{d}$ describes the domain with boundary $\partial\Omega$.
-PDEs often arise in high-dimensional settings (i.e., $d$ is large); for example, well-known instances such as the Poisson equation (Yu et al., 2018), the linear conservation law (Chen et al., 2021), and the nonlinear Schrödinger equation (Han et al., 2020) regularly involve multiple spatial and temporal variables.
-While low-dimensional PDEs can be solved reliably and efficiently via traditional mesh-dependent methods such as finite difference methods (LeVeque, 2007), finite elements methods (Hughes, 2003), finite volume methods (Toro, 2013), and spectral methods (Boyd, 2001), developing efficient and accurate algorithmic frameworks for computing numerical solutions to high-dimensional PDEs remains an important and challenging topic due to the curse of dimensionality (i.e., the computational costs grow exponentially with respect to the dimensionality of the problem’s domain) (E et al., 2021).
-Before solving a PDE, one must understand certain fundamental properties of the solution $u(x)$ that follow from the PDE itself and from the data—namely, $f(u(x),x)$ and $g(x)$.
-These considerations align with the classical theory of PDEs (Evans, 2022), which focuses on establishing well-posedness (i.e., the existence and uniqueness of solutions) and their regularity and continuity.
+$Du=f,\quad\text{in }\Omega,\quad Bu=g,\quad\text{on }\partial\Omega,$ (1)
+where $D$ denotes a differential operator defining the dynamics, $Bu=g$ represents the boundary condition (e.g., Dirichlet, Neumann, and Robin), and $\Omega\subseteq\mathbb{R}^{d}$ is the domain with boundary $\partial\Omega$.
+PDEs often arise in high-dimensional settings (i.e., $d$ is large); for example, well-known instances such as the Poisson equation [43], the linear conservation law [5], and the nonlinear Schrödinger equation [17] naturally involve multiple spatial and temporal variables.
+While low-dimensional PDEs can be solved effectively via traditional mesh-dependent methods such as finite difference methods [26], finite elements methods [21], finite volume methods [42], and spectral methods [3], developing efficient and accurate algorithmic frameworks for computing numerical solutions to high-dimensional PDEs remains an important and challenging topic due to the curse of dimensionality (i.e., the computational costs grow exponentially with respect to the dimensionality) [10].
+Before solving a PDE, one must understand certain fundamental properties of the solution $u$ that follow from the PDE itself and from the data, namely, $f$ and $g$.
+These considerations align with the classical theory of PDEs [11], which focuses on establishing well-posedness (i.e., the existence and uniqueness of solutions) and their regularity and continuity.
 However, the underlying symbolic relationship among $u$, $f$ and $g$ often remains underexplored.
 To address this gap, we pose the following question:
-Assuming that the PDE (1) admits an analytical solution $u$, and given the operators appearing in the expressions of $f$ and $g$, what are the corresponding operators in the expression for $u$?
+Assuming that the PDE (1) admits an analytical solution $u$, and given the operators in $f$ and $g$, how are the operators in $u$ related to those in $f$ and $g$?
 Understanding this symbolic connection is crucial for deriving fully interpretable, closed-form solutions, a feature lacking in most existing solution methods, particularly in the high-dimensional setting.
-As an illustrative example, we consider solving the PDE (1) via the least square method (Dissanayake and Phan-Thien, 1994; Lagaris et al., 1998; Sirignano and Spiliopoulos, 2018), which defines a straightforward functional to characterize the error of the estimated solution by
+As an illustrative example, we consider solving the PDE (1) via the least square method [9, 24, 37],which defines a straightforward loss to characterize the error of the estimated solution by
 $L(u)=\|Du-f\|^{2}_{L^{2}(\Omega)}+\lambda\|Bu-g\|^{2}_{L^{2}(\partial\Omega)},$ (2)
 where $\lambda>0$ balances the influence of boundary conditions. The goal is to find a function $u^{*}\in L^{2}(\Omega)$ so that the least square loss is minimized, i.e.,
-$u^{*}=\mathrm{argmin}_{u\in L^{2}(\Omega)}\;L(u).$ (3)
+$u^{*}=\mathrm{argmin}\left\{L(u)\;:\;u\in L^{2}(\Omega)\right\}.$ (3)
 Because the search space $L^{2}(\Omega)$ is an infinite dimensional space, finding the solution $u^{*}$ is extremely challenging.
 Fortunately, the search space can be significantly simplified if one knows the operators appearing in the solution in advance, making the search of the optimal solution much easier and effective.
-This can be accomplished by adopting the recently developed finite expression method (FEX) (Liang and Yang, 2022).
-To uncover the symbolic relationship described above, we leverage large language models (LLMs) as predictive tools.
+This can be accomplished by adopting state-of-the-art symbolic machine learning approaches, including the efficient finite expression method (FEX) of [30].
+To uncover the symbolic relationship, we leverage large language models (LLMs) as predictive tools.
 The process begins with the generation of a comprehensive, structured dataset composed of symbolic expressions derived from a diverse range of PDE types, including elliptic, parabolic, and hyperbolic equations, along with their associated boundary and initial conditions.
@@
 1.1 Related work
-Artificial intelligence (AI) for Science takes advantage of modern advanced machine learning (ML) methods to facilitate, accelerate and enhance scientific discovery (Stevens et al., 2020).
+Artificial intelligence (AI) for Science takes advantage of modern advanced machine learning (ML) methods to facilitate, accelerate and enhance scientific discovery [40].
 Its importance lies in the ability to handle massive datasets, uncover hidden patterns, and simulate complex phenomena.
-Inspired by the remarkable success of AI across various domains, deep neural network (DNN)-based approaches (LeCun et al., 2015; Goodfellow, 2016) have become increasingly popular for solving (high-dimensional) PDEs.
+Inspired by the remarkable success of AI across various domains, deep neural network (DNN)-based approaches [25, 13] have become increasingly popular for solving (high-dimensional) PDEs.
 These methods offer certain appealing advantages over traditional numerical techniques, making them a compelling alternative for tackling complex problems.
-First, DNN-based approaches empirically mitigate the curse of dimensionality that plagues conventional discretization approaches, leading to more efficiency in handling high-dimensional PDEs Han et al.
-(2018).
-Second, DNNs, acting as mesh-free universal function approximators (Cybenko, 1989; Hornik et al., 1989; Goodfellow, 2016), can be trained to learn highly complex PDE solutions without the need of prespecifying basis functions or meshes (Berg and Nyström, 2018; Raissi et al., 2019).
+First, DNN-based approaches empirically mitigate the curse of dimensionality that plagues conventional discretization approaches, leading to more efficiency in handling high-dimensional PDEs [16].
+Second, DNNs, acting as mesh-free universal function approximators [7, 19, 13], can be trained to learn highly complex PDE solutions without the need of prespecifying basis functions or meshes [1, 33].
 Hence, they are highly suitable to approximate complex solution spaces that challenges the traditional mesh-based approaches.
-Third, the solutions represented by DNNs can be evaluated efficiently once trained, which turns out to be a crucial property for tasks, including uncertainty quantification and real-time optimal control, require calling PDE solutions repeatedly (Zhu et al., 2019).
-Last but not least, continuous progress in designing comprehensive neural network architectures—such as Fourier neural operators and advanced physics-informed neural network frameworks—alongside theoretical advancements in understanding their approximation capabilities, are making these methods increasingly applicable (Lu et al., 2019; Li et al., 2020).
+Third, the solutions represented by DNNs can be evaluated efficiently once trained, which turns out to be a crucial property for tasks, including uncertainty quantification and real-time optimal control, require calling PDE solutions repeatedly [46].
+Last but not least, continuous progress in designing comprehensive neural network architectures, such as Fourier neural operators and advanced physics-informed neural network frameworks, alongside theoretical advancements in understanding their approximation capabilities, are making these methods increasingly applicable [31, 28].
 However, several critical factors often hinder deep neural networks (DNNs) from achieving highly accurate solutions, even for relatively simple problems.
-These include the need for large, diverse, and high-quality training data (Szegedy, 2013; Buda et al., 2018), sensitivity to hyperparameter selection (Bergstra and Bengio, 2012), the challenge of optimizing highly nonconvex objectives (Dauphin et al., 2014; Choromanska et al., 2015), issues like vanishing and exploding gradients (Glorot and Bengio, 2010), and poor generalization capabilities (Zhang et al., 2021).
-Additionally, solutions produced by DNNs often lack interpretability, preventing users from leveraging insights for future decision-making (Rudin, 2019).
-To overcome these challenges and achieve both highly accurate and interpretable solutions for PDEs, Liang and Yang (2022) recently introduced the finite expression method (FEX).
-This approach leverages recent advances in approximation theory, which demonstrate that high-dimensional functions can be effectively approximated using functions composed of simple operators (Shen et al., 2021a, b; Zhang et al., 2022).
+These include the need for large, diverse, and high-quality training data [41, 4], sensitivity to hyperparameter selection [2], the challenge of optimizing highly nonconvex objectives [8, 6], issues like vanishing and exploding gradients [12], and poor generalization capabilities [44].
+Additionally, solutions produced by DNNs often lack interpretability, preventing users from leveraging insights for future decision-making [34].
+To overcome these challenges and achieve both highly accurate and interpretable solutions for PDEs, [30] recently introduced the finite expression method (FEX).
+This approach leverages recent advances in approximation theory, which demonstrate that high-dimensional functions can be effectively approximated using functions composed of simple operators [35, 36, 45].
 This methodology seeks approximate PDE solutions within the space of functions composed of finitely many analytic expressions, effectively avoiding the curse of dimensionality.
@@
 This leads to a combinatorial optimization (CO) problem, which is then solved by a deep reinforcement learning (RL) method with the policy gradient optimizer.
-The FEX outperforms certain existing numerical PDE solvers in terms of accuracy, interpretability and memory efficiency, making it a promising approach for solving high-dimensional PDEs and other complex problems; see e.g., (Song et al., 2023, 2024; Hardwick et al., 2024).
+The FEX outperforms certain existing numerical PDE solvers in terms of accuracy, interpretability and memory efficiency, making it a promising approach for solving high-dimensional PDEs and other complex problems; see e.g., [38, 39, 18, 29].
+2 Theoretical Insight for Poisson Equations
+While providing a complete answer to the question raised in the Introduction remains a significant challenge, in this section we take an initial step toward addressing it by offering theoretical insights that illuminate key aspects of the problem and motivate directions for future research.
+Specifically, we focus on the classical Poisson equation with Dirichlet boundary condition
+$-\Delta u=f\quad\text{in }\Omega,\qquad u=g\quad\text{on }\partial\Omega,$ (4)
+where $\Omega\subseteq\mathbb{R}^{d}$ is a compact domain.
+By analyzing this simple class of elliptic PDEs under, we aim to shed light on the structural properties and solution behavior relevant to the broader question, thereby laying the groundwork for more comprehensive theoretical development in future work.
+Theorem 1.
+Let $\Omega\subset\mathbb{R}^{d}$ be a compact domain with boundary $\partial\Omega$ such that the distance function $\mathcal{D}(x):=\mathrm{dist}(x,\partial\Omega)$ admits an explicit analytical expression.
+Suppose the source term $f\in C^{0,\alpha}(\Omega)$ for some $\alpha\in(0,1]$ (i.e., $f$ is Hölder continuous), and the boundary condition $g$ is the restriction to $\partial\Omega$ of a function in $C^{2}(\overline{\Omega})$.
+Let $u$ be the analytical solution to the Possion equation (4).
+Then for any $\delta\in(0,1)$, there exists an approximating function $\tilde{u}:\mathbb{R}^{d}\to\mathbb{R}$ such that:
+  1. 1.
+    $\|u-\tilde{u}\|_{L^{2}(\Omega)}\leq\delta$,
+  2. 2.
+    $\tilde{u}$ is constructed using only the elementary arithmetic operators +,×$+,\;\times,\;(\cdot)^{2}$ and the operators appearing in the evaluation of $\mathcal{D}(\cdot)$, $f(\cdot)$, and $g(\cdot)$,
+  3. 3.
+    The total number of operations needed to evaluate $\tilde{u}(x)$ at any $x\in\mathbb{R}^{d}$ satisfies:
+    $\texttt{ops}(\tilde{u})=\texttt{poly}(\delta^{-1})\cdot\left(\texttt{ops}(f)+\texttt{ops}(g)\right),$
+    where $\texttt{poly}(\delta^{-1})$ denotes a quantity polynomial in $\delta^{-1}$.
+Proof.
+See Appendix B. ∎
+This theorem offers a partial constructive resolution to the open research question posed in the previous section: it identifies the specific operators needed to represent a $\delta$-accurate approximation of the PDE solution $u$ in terms of the operators used to define the source term $f$, the boundary condition $g$, and the domain geometry via the distance function $\mathcal{D}$.
+The central insight is that an accurate approximation $\tilde{u}$ can be constructed using only a limited set of algebraic operations, along with the computational building blocks inherent in $f$, $g$, and $\mathcal{D}$.
+This result carries several important implications.
+(1) In contexts where one seeks symbolic models for physical phenomena, the theorem guarantees the existence of a low-complexity symbolic approximation of $u$ that relies solely on a structured and interpretable set of primitive operations.
+(2) It also informs the design of neural architectures for PDE approximation, suggesting that imposing operator priors, i.e., by reusing the operators present in $f$ and $g$, is sufficient to achieve arbitrarily high approximation accuracy, thereby promoting efficiency and interpretability.
+(3) The polynomial bound on computational complexity with respect to $\delta^{-1}$ underscores the practical feasibility and scalability of such approximations.
+Taken together, these insights lay a foundation for a broader theory of operator-preserving approximations for general PDEs, with significant potential impact in scientific machine learning, computational physics, and algorithmic model design.
 3 Methodology
@@
 Below, we detail the pipeline for generating synthetic data, the procedure for fine-tuning the LLMs, and the evaluation of their predictive capabilities.
-2.1 Binary Computational Trees for PDE Expressions
-Expressions in PDEs are represented as binary computational trees, as depicted in Figure 1.
+3.1 Binary Computational Trees for Expressions
+Expressions in PDEs are represented as binary computational trees, as depicted in Figure 5 in Appendix A.
 These trees provide a structured and hierarchical way to encode mathematical expressions, enabling efficient exploration and manipulation of the solution space.
-The tree generation process is systematic and begins by specifying the depth of the binary tree, which determines the complexity and granularity of the resulting expressions.
-Figure 1: Computational expression tree (Liang and Yang, 2022).
-The construction of the tree involves the following steps:
-  1. 1.
-    Depth Specification: The depth of the tree is predefined, controlling the number of hierarchical levels in the tree.
-A deeper tree allows for more complex expressions but increases computational cost.
-  2. 2.
-    Recursive Construction: At each level of depth, the tree is built recursively.
-Starting from the root node, each node is expanded by adding child nodes according to the rules defined for operators.
-This recursive process ensures a systematic exploration of the solution space, covering a wide range of possible expressions.
-Each node in the computational tree represents an operator, which can be either a unary operator or a binary operator:
-  • Unary Operators: These operators act on a single operand and include functions such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other nonlinear transformations ($\sqrt{\cdot}$, ||$\cdot$|| ⋅ |, $(\cdot)^{k}$, $k\in\mathbb{Z}$).
-For each node with a unary operator, two scalars ($\alpha$ and $\beta$) are coupled with the node.
-These scalars enable the composition of the underlying operation with an affine function, resulting in expressions of the form $\alpha\cdot u(\cdot)+\beta$, where $u(\cdot)$ is the unary operation.
-This coupling ensures that the tree can represent both nonlinear transformations and linear adjustments, enhancing its expressive power.
-  • Binary Operators: These operators require two operands and include standard arithmetic operations including addition ($+$), subtraction (--), multiplication ($\times$) and division (///).
-Binary operators combine two sub-expressions (left and right child nodes) to form more complex expressions.
-For example, a multiplication node might combine the results of two subtrees representing different terms in an expression.
-The use of binary computational trees offers several advantages: (1) The tree structure provides a clear and interpretable representation of mathematical expressions, making it easier to analyze and manipulate them; (2) By recursively constructing trees of varying depths, the method ensures a thorough exploration of the solution space, capturing both simple and complex expressions; and (3) The inclusion of both unary and binary operators, along with affine transformations, allows the tree to represent a wide range of mathematical operations and relationships.
-By systematically generating and evaluating computational trees, it becomes possible to identify candidate expressions that best describe the underlying dynamics of a system.
-Additionally, the framework can be combined with advanced techniques, including optimization and reinforcement learning, to refine the scalars and operators, further improving the accuracy of the discovered models.
+The construction of the tree proceeds in two main stages.
+Depth Specification: The maximum depth of the tree is specified in advance, determining the number of hierarchical levels and thereby controlling the expressive capacity of the resulting representations.
+While deeper trees enable the generation of more complex expressions, they also entail higher computational costs.
+Recursive Construction: Given the specified depth, the tree is constructed recursively.
+Starting from the root node, each node is expanded by generating child nodes according to predefined operator rules.
+Each node in the computational tree encodes an operator, which may be either a unary or binary operator.
+Unary Operators: Unary operators act on a single operand and include a variety of nonlinear transformations such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other elementary operations (e.g., $\sqrt{\cdot}$, ||$\cdot$|| ⋅ |, and power functions $(\cdot)^{k}$ for $k\in\mathbb{Z}$).
+Each unary operator node is augmented with two scalar parameters, $\alpha$ and $\beta$, which enable the composition of the operator with an affine transformation, yielding expressions of the form $\alpha\cdot u(\cdot)+\beta$, where $u(\cdot)$ denotes the unary operation.
+This parameterization allows the tree to flexibly represent both nonlinear and linearly adjusted transformations, thereby enhancing its expressiveness.
+Binary Operators: Binary operators operate on two operands and include the standard arithmetic operations: addition ($+$), subtraction (--), multiplication ($\times$), and division (///).
+Each binary operator node combines the outputs of its left and right child subtrees to produce more complex expressions.
+For example, a multiplication node may combine two distinct sub-expressions, thereby enabling the construction of richer symbolic forms through hierarchical composition.
 3.2 Data Generation for Equation Types
-In this subsection, we describe the data generation process for constructing a structured dataset of symbolic expressions derived from various PDE types.
-The dataset is designed to capture the symbolic relationships between the solution of a PDE and the associated problem data, including the boundary conditions and right-hand-side functions.
-The data generation process is outlined in Figure 2.
+We next describe the data generation process for constructing a structured dataset of symbolic expressions derived from various PDE types. The data generation process is outlined in Figure 1.
 Figure 1: Data generation pipeline.
-The construction of the one data point involves the following steps.
-PDE Type Specification.
-We first specify the type of PDE (e.g., Poisson, linear conservation law, and Schrödinger) and the type of boundary conditions (e.g., Dirichlet, Neumann and Robin) based on the information of the operators $D$ and $B$.
-This results in the first component of the data point, which includes the PDE type and boundary condition type of the data point.
-Randomly generated function $u(x)$.
-Next, we generate a random function $u(x)$ by constructing a binary computational tree with a specified depth, as described in Section 2.1.
-The tree is populated with random unary and binary operators drawn from prescribed operator sets that are rich enough.
-Consequently, we obtain the expression tree for the solution function $u(x)$, which is used to derive the second component of the data point, i.e., expressions of the right-hand-side function $f$ and the function $g$ defined on the boundary of the domain.
+Specifically, the construction of a single data point consists of the following steps: PDE Type Specification.
+We begin by specifying the type of PDE and the type of boundary condition, based on the associated differential and boundary operators $D$ and $B$.
+This information forms the first component of the data point, identifying the PDE and boundary condition types.
+Generation of Random Function $u$.
+Next, we construct a random solution function $u$ using a binary computational tree of predefined depth, as described in Section 3.1.
+The tree is populated with randomly selected unary and binary operators from a sufficiently expressive operator set.
+This yields a symbolic expression for $u$, from which we compute the corresponding right-hand-side function $f$ and boundary function $g$, together forming the second component of the data point.
 Postfix Representation.
-After obtaining the expressions of the functions $u(x)$, $f(u(x),x)$, and $g(x)$, we convert them into postfix notation.
-This conversion simplifies the tokenization process and aligns the expressions with the input format required by LLMs.
-The postfix notation also eliminates the need for parentheses, making the expressions more interpretable and easier to parse.
-More importantly, it reflects the relative positional relationship of operators in the expression, which is crucial for the LLMs to learn the symbolic relationships between the operators.
-In the end, we obtain the second component of the data point, which includes the operators used in the right-hand-side function $f(u(x),x)$, and the function $g(x)$.
-These input component are then concatenated to generate the input.
-Meanwhile, the label of the data set is obtained similarly by converting the expression of the function $u(x)$ into postfix notation.
-See Figure 3 for the detailed convertion pipeline and a simple illustrated example.
-Figure 3: Pipeline for converting data into postfix notation and an illustrated example.
-2.3 Fine-tuning
-Inspired by the success of LLMs, we propose fine-tuning the sequence-to-sequence Transformer architecture that leverages cross-attention to incorporate PDE context when generating solution operators.
-This architecture is selected because it effectively captures long-range dependencies and the complex hierarchical structure inherent in PDE metadata, thereby learning an effective mapping between symbolic operator sequences.
-We first preprocess and tokenize the dataset generated as in the previous subsection, and then fine-tune the LLM using the proceed data using the common practices in the literature.
-The whole pipeline is outlined in Figure 4 and the key components in the pipeline are discussed in detailed as follows.
-Figure 4: Pipeline for tokenization and embedding of operator sequences for LLM fine-tuning
+To facilitate processing by large language models (LLMs), the expressions for $u$, $f$, and $g$ are converted into postfix notation.
+This format simplifies tokenization, avoids the use of parentheses, and explicitly encodes operator precedence through position.
+Importantly, it captures the structural relationships among operators, which helps LLMs learn symbolic patterns.
+The postfix expressions of $f$ and $g$ constitute the model input, while the postfix form of $u(x)$ serves as the output label.
+3.3 Fine-tuning LLM
+With a generated data set, we can then fine-tune a certain LLM to learn the symbolic relastion in PDEs.
+The whole pipeline is summarized in Figure 2, and we will describe each component in detail for the rest of this subsection.
+Figure 2: Overview of the fine-tuning pipeline
+Model Selection.
+Encoder-decoder and decoder-only architectures are two fundamental types of LLMs, each with distinct design principles and strengths.
+Encoder-decoder models (e.g., T5 [32] and BART [27]) consist of an encoder that transforms the entire input sequence into a rich latent representation, and a decoder that generates the output sequence token by token, attending to the encoder’s outputs via cross-attention.
+This architecture is well-suited for tasks such as machine translation, summarization, and question answering, where understanding the full input before generating the output is crucial.
+In contrast, decoder-only models (e.g., Llama3 [14]) treat the input and output as a single continuous sequence and generate tokens autoregressively using causal self-attention, meaning each token can only attend to previous tokens.
+This design simplifies the model architecture and training process, and has proven particularly effective for open-ended context generation and few-shot prompting.
+Based on our empirical results, decoder-only models are preferable for our task, as they more accurately predict the symbolic operators associated with PDE solutions.
 Tokenization.
-We begin by loading the PDE data in jsonl format, shuffling it, and splitting it into training and validation sets.
-Each data point is then processed by a textual embedding layer that converts operator symbols (e.g., sin, cos, exp) into tokens.
-To ensure a robust and consistent representation, we use HuggingFace’s  [^1] https://github.com/huggingface AutoTokenizer, which selects an appropriate tokenization scheme based on the underlying pre-trained model.
-The tokenizer pads sequences to a fixed length and incorporates learnable positional embeddings.
-Additionally, an attention mask is generated to differentiate between valid and padding tokens, facilitating self-attention computations.
-Finally, padding tokens in the label sequences are replaced with $-100$, the default ignore_index in PyTorch’s  [^2] https://pytorch.org/ nn.CrossEntropyLoss, ensuring they are ignored during loss computation.
-This approach standardizes the transformation of textual operator representations into numerical tokens for model training.
-Architecture and Finetuning: Next, we select a pretrained LLM with a sequence-to-sequence Transformer architecture.
-We fine-tune this model using the HuggingFace Trainer API, which provides a suite of tools for efficient optimization.
-During fine-tuning, we minimize the token-wise cross-entropy loss, which penalizes incorrect token predictions, including misclassification of the End-of-Sequence (EOS) token.
-Backpropagation is performed to update the model’s parameters according to the optimization algorithm.
-We emphasize the use of teacher forcing  [^3] Teacher forcing is a training strategy where, at each time step, the decoder is provided with the actual ground-truth token instead of its own previous prediction.
-during fine-tuning.
-By supplying the decoder with the full target sequence, this approach ensures that the token-wise cross-entropy loss is computed over aligned positions, preventing error accumulation and reinforcing the correct sequential dependencies more effectively and robustly.
+We next perform tokenization to convert raw text into tokens that the model can process numerically, enabling efficient representation and learning.
+Specifically, each data point (in raw text) is serialized into a prompt followed by its target operator sequence, as demonstrated in the following.
+$\underbrace{\texttt{Type:}<\mathrm{PDE}>\,\texttt{|}\,\texttt{RHS:}<\mathrm{ops}>\,\texttt{|}\,\texttt{BC\_Type:}<\mathrm{bc}>\,\texttt{|}\,\texttt{Solution:}}_{\text{PROMPT}}\;\|\|\underbrace{\texttt{<op}_{1}\ \dots\ \texttt{op}_{n}\,\texttt{<EOS>}\texttt{>}}_{\text{TARGET}}$
+Next, a pretrained tokenizer is applied to convert the input text into its corresponding token sequence.
+In this work, we use the SentencePiece  [^1] Available at https://github.com/google/sentencepiece.
+tokenizer and explicitly set the PAD token to match the EOS token, ensuring that padding positions are never treated as valid outputs.
+To accommodate memory constraints, we truncate the prompt to a maximum of 256 tokens and the target to a maximum of 64 tokens, resulting in a concatenated sequence length of 300.
+Additionally, we construct a binary attention mask, where real tokens are marked with 1 and padding positions with 0.
+This mask is supplied to the model to ensure that self-attention computations ignore padded positions.
+Fine-tuning.
+Given a pretrained decoder-only language model and a tokenized dataset, we proceed to the fine-tuning phase, where the model’s parameters are updated via stochastic gradient-based optimization.
+Specifically, we minimize the cross-entropy loss between the model’s predicted token distributions and the ground truth tokens in the training data.
+This objective encourages the model to assign higher probabilities to correct next-token predictions, thereby adapting its knowledge to the target task or domain.
+For optimization, we use the AdamW optimizer, a variant of Adam [22] that decouples weight decay from the gradient updates, which has been shown to improve generalization performance in transformer-based models.
+To further enhance the efficiency of training, we adopt several computational techniques.
+First, we employ mixed-precision training, which allows computations to be performed in lower precision (e.g., float16) while maintaining a decent level of stability and accuracy.
+This significantly reduces memory usage and accelerates training.
+Second, we leverage Low-Rank Adaptation (LoRA) [20], a parameter-efficient fine-tuning technique that injects trainable low-rank matrices into specific layers (typically attention or feedforward layers) of the pretrained model while keeping the original model weights frozen.
+This approach greatly reduces the number of trainable parameters and memory footprint during training, making it especially suitable for large-scale models and resource-constrained environments.
+Together, these strategies enable effective and efficient fine-tuning of large language models on our tasks.
 3.4 Inferencing
-Once trained, the model can predict operator sets in solutions for new PDEs.
-Specifically, we provide the language model with a data point consisting of the PDE type, its right-hand side (RHS), and the associated boundary condition, as described in Section 2.2.
-This data is then tokenized following the procedure outlined in the previous subsection.
-The language model’s decoder then auto-regressively generates an operator sequence representing the solution, with each predicted token conditioned on previously decoded tokens and the input context.
-While the exact sequential order may vary slightly, the model consistently identifies the correct set of operators, prioritizing presence over strict ordering.
-Since the model’s output is a numerical vector, a final post-processing step is required to extract the predicted operators.
-First, we extract the unique operators by removing duplicates from the predicted postfix representation, ensuring that only essential components remain.
-Because the model is trained to reproduce entire operator sequences rather than direct sets, this step distills the output into a concise and meaningful form.
-Next, the obtained set is checked for misspellings and tokenization errors.
-Given that the model generates sequences based on learned patterns rather than strict syntax rules, occasional character-level errors may arise, particularly with uncommon symbols or operators.
-To ensure correctness and consistency, any malformed operators are discarded.
-To evaluate the accuracy of the fine-tuned LLM, we use the squared norm to quantify the discrepancy between the predicted operator set and the ground-truth label.
-Particularly, given a dictionary of all possible operators of size $n$, each operator set is represented as a binary vector, as illustrated in Table 1.
-Binary Vector
-Function Preprocessed Operator Set
-$x_{1}$
-$x_{2}$
-^2
-^3
-SIN
-COS
-EXP
-$(5x_{1})^{2}+\sin(3x_{2})\cdot x_{2}$ [ $x_{1}$, $x_{2}$, ^2, +, *, SIN ] 1 1 1 0 1 1 1 0 0
-$5\exp(2x_{1})\cdot(\cos(6x_{1}))^{3}$ [ $x_{1}$, ^3, *, COS, EXP ] 1 0 0 1 0 1 0 1 1
-Table 1: Sample binary-vector encoding for two example expressions. The size of the dictionary is 9.
-Then, the squared distance between two such operator sets, represented by vectors $y\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{n}$, can be defined as
-$\|y-z\|^{2}=\sum_{i=1}^{n}\left(y_{i}-z_{i}\right)^{2}.$
-Clearly, it measures the number of mismatched operators between the two operator sets.
-Moreover, this metric prioritizes the presence of operators over their sequence order, ensuring that the predicted operators remain useful for numerical and analytical PDE solving, as well as for scientific analysis of PDE models.
+Once fine-tuned, the language model is capable of predicting the postfix representation of previously unseen partial differential equation (PDE) solutions.
+Given a new input instance, comprising the PDE type, its right-hand side, and the corresponding boundary conditions, the input is first tokenized and then fed into the model.
+The decoder will auto-regressively generate a sequence of tokens that represents the solution in postfix notation, encoding the sequence of mathematical operators and operands that define the PDE solution.
+Since the model’s raw output is a token sequence, a post-processing pipeline is required to extract a clean and interpretable set of predicted operators.
+The first step involves identifying the unique set of operators by removing duplicate tokens from the generated sequence.
+Following this, we perform error checking to detect and eliminate misspelled or malformed tokens that may result from tokenization artifacts or character-level prediction errors (issues that can arise because the model learns statistical patterns rather than enforcing strict syntactic correctness).
+Such errors are particularly likely when dealing with rare or domain-specific mathematical symbols.
+Any invalid or unrecognized tokens are subsequently discarded to ensure the integrity of the final operator set.
+To quantitatively evaluate the model’s predictive accuracy, we compute the squared $\ell_{2}$ norm of the difference between the predicted and ground-truth operator sets.
+Specifically, we encode each operator set as a binary vector over a fixed dictionary of $n$ possible operators, where each vector component indicates the presence or absence of a given operator.
+This encoding transforms the comparison into a well-defined vector space problem.
+An illustrative example of this binary encoding scheme can be found in Appendix C.1.
+Then, the squared distance between two such operator sets, represented by two binary vectors $y\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{n}$, can be defined as $\|y-z\|^{2}=\sum_{i=1}^{n}\left(y_{i}-z_{i}\right)^{2}.$ Clearly, it measures the number of mismatched operators between the two operator sets.
 4 Experimental Results
@@
 4.1 Experimental Setup
-For data generation, we consider two types of PDEs—namely, the Poisson equation and the linear conservation law—each paired with three commonly used boundary conditions: Cauchy, Dirichlet, and Neumann.
-Using a tree depth of 3, we randomly generate a dataset of 60,000 equations, with 30,000 examples for each PDE type, as described in Section 2.2.
-This dataset is then used to fine-tune two state-of-the-art large language models: BART (Lewis et al., 2019) and T5 (Raffel et al., 2020).
-For fine-tuning, we optimize the cross-entropy loss using the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and a weight decay of $0.01$.
-The models are trained for 3 epochs with a batch size of 2.
-Fine-tuning is conducted on an NVIDIA A6000 GPU with 48 GB of memory.
-To efficiently manage storage, we save checkpoints every 500 steps and retain only the three most recent checkpoints.
-Finally, we integrate the model into the FEX solver [^4] For a detailed description of the FEX method, we refer the reader to Liang and Yang (2022), and for the implementation of the original FEX, see https://github.com/LeungSamWai/Finite-expression-method..
-During the execution of FEX for a given PDE, the model is invoked for inference.
-After inference, the post-processing step detailed in Section 2.4 is applied to extract the unique unary and binary operator sets.
-These sets are then used to configure the available operators within FEX during its search for an analytical PDE solution.
-We refer to this version of FEX as the "LLM-informed FEX" and denote the original FEX as "uninformed FEX." The uninformed FEX employs a fixed operator set:
+For data generation, we consider the Poisson equation and the Linear Conservation Law, each paired with three commonly used boundary conditions: Cauchy, Dirichlet, and Neumann conditions.
+Using a tree depth of 3, we randomly generate a dataset of 198,000 equations, with 99,000 examples for each PDE type.
+This dataset is then used to fine-tune 4 state-of-the-art large language models: BART [27], T5 [32], LLaMA-3B and LLaMA-8B [14].
+For fine-tuning, we minimize the cross-entropy loss using the AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay coefficient of $0.01$.
+Models are trained for 12 epochs with a batch size of 32.
+Training is performed on 8 NVIDIA A6000 GPUs, each equipped with 48 GB of memory.
+To improve training efficiency and scalability, we employ mixed-precision training (BF16) and gradient checkpointing.
+We also apply Low-Rank Adaptation (LoRA) with a rank of 16, scaling factor of $32$, and dropout rate of 0.1.
+Model evaluation and checkpointing are conducted at the end of each epoch.
+Finally, we integrate the model into the FEX solver.
+[^2] For a detailed description of the FEX method, see [30]; for the original implementation, visit https://github.com/LeungSamWai/Finite-expression-method.
+During the execution of FEX on a given PDE, the fine-tuned model is called for inference.
+The resulting expression is post-processed to extract the sets of unique unary and binary operators, which are then used to define the operator pool available to FEX during its symbolic search for an analytical solution.
+We refer to this enhanced version as LLM-informed FEX, in contrast to the original uninformed FEX.
+The uninformed FEX employs a fixed operator set:
 […]
@@
 4.2 Effectiveness of Fine-tuned LLMs
-In this subsection, we evaluate the performance of the fine-tuned BART and T5 models in predicting operator sets for the two PDE types.
-Specifically, we track the training and validation losses throughout the training process, as well as the average number of mismatched operators on the test dataset for each epoch.
-The computational results are summarized in Tables 2 and 3.
-Epoch Training Loss Validation Loss Average Mismatch
-1 0.058200 0.057948 2.475833
-2 0.040000 0.046193 2.020000
-3 0.039200 0.038100 1.720210
-Table 2: Fine-tuning BART on large dataset.
-Epoch Training Loss Validation Loss Average Mismatch
-1 0.046800 0.045303 1.711056
-2 0.037200 0.036140 1.544056
-3 0.012300 0.020392 1.012000
-Table 3: Fine-tuning T5 on large dataset.
-Both fine-tuned models exhibit a consistent decrease in training and validation loss across epochs, indicating successful fine-tuning.
-However, T5 demonstrates a more pronounced reduction in both loss metrics.
-The larger gap between BART’s training and validation loss suggests that it may be more prone to overfitting compared to T5, which appears to generalize better.
-In terms of the average number of mismatched operators, T5 consistently outperforms BART in this regard across all epochs.
-The sharper decline in mismatch for T5 suggests that it adapts more effectively to the symbolic representation of PDE solutions and captures the underlying operator structures with higher precision.
-Overall, T5 demonstrates superior performance in predicting operators for PDE solutions, making it a more suitable choice for integration into the LLM-informed FEX solver.
-In light of these findings, we incorporate the fine-tuned T5 model into the FEX framework to enhance its operator selection process.
-3.3 LLM-informed FEX
-In this subsection, we evaluate the effectiveness of the LLM-informed FEX using the fine-tuned T5 model in comparison to the uninformed FEX.
-As an initial validation step, we conduct a sanity check using several simple examples to ensure the correctness of the approach.
-To comprehensively assess the performance of the proposed framework, we further test both versions of FEX on a dataset of 100 randomly generated PDEs, consisting of 50 Poisson equations and 50 linear conservation law equations.
-3.3.1 Sanity Check
-We first construct 4 Poisson equations and 4 linear conservation law equations, whose true solutions, denoted as $u(x)$, are randomly handcrafted.
-For each example, we present the Binary and Unary set sizes, the number of iterations and computational time taken to find the approximate solution, and the error with respect to the true solution.
-The results are summarized in Tables 4 and 5, respectively.
-$u(x)$ Method Binary Size Unary Size Iters Time [m] Error
-$4\,\cos\bigl{(}4x^{2}\cos(x_{0})\bigr{)}$ LLM-informed 1 2 4.25 8.25 0
-Uninformed 3 9 167 340 0
-$4\,x_{1}^{3}+4\,x_{1}^{2}+2\,\cos\bigl{(}4\,x_{1}^{3}\cos(x_{0})\bigr{)}$ LLM-informed 2 4 102 286 $10^{-8}$
-Uninformed [^5] Uninformed FEX did not converge within 2000 iterations, indicated by “2000+” (iterations) and “N/A” for error. 3 9 2000+ 2400+ N/A
-$128\,x_{2}^{3}+2\,x_{2}\,e^{4\,x_{2}^{4}}$ LLM-informed 2 4 34.5 40.5 $4\times 10^{-7}$
-Uninformed 3 9 90 186 $6\times 10^{-7}$
-$64\,x_{1}^{2}\,e^{2\,x_{2}}$ LLM-informed 1 3 15.5 21 $3\times 10^{-8}$
-Uninformed 3 9 103.5 161 $3\times 10^{-8}$
-Table 4: Sanity check for the Poisson equation.
-$u(x)$ Method Binary Size Unary Size Iters Time [m] Error
-$\sin\Bigl{(}x_{1}+\tfrac{\pi}{4}x_{2}+\tfrac{\pi}{4}x_{3}\Bigr{)}$ LLM-informed 2 2 1 0.3 0
-Uninformed 3 9 3 1 0
-$4\,\bigl{(}2x_{0}-2x_{2}^{3}\bigr{)}^{3}$ LLM-informed 2 2 2 0.6 0
-Uninformed 3 9 33.3 9.4 0
-$8\,x_{2}^{2}+8\,\sin(x_{1})$ LLM-informed 2 3 1 1 0
-Uninformed 3 9 24 28.5 0
-[…]
-Uninformed 3 9 273 488 0
-Table 5: Sanity check for the linear conservation law euqation.
-The computational results presented in Tables 4 and 5 demonstrate that the LLM-informed FEX significantly improves efficiency over the uninformed FEX while maintaining high accuracy.
-One major advantage is the reduction in the number of required operators.
-The uninformed FEX consistently uses a fixed operator set of 3 binary and 9 unary operators, whereas the LLM-informed FEX dynamically selects a smaller, more relevant subset—typically with at most 2 binary and 4 unary operators.
-This reduced search space leads to a dramatic decrease in the number of iterations required for convergence.
-For instance, in the first Poisson equation example, the LLM-informed FEX requires only 4.25 iterations compared to 167 for the uninformed version, and in another case, the uninformed FEX exceeds 2000 iterations without converging, while the LLM-informed approach successfully completes the solution in 102 iterations.
-This efficiency gain translates into significant computational time savings, with the LLM-informed approach solving problems up to an order of magnitude faster.
-Despite this drastic reduction in computational cost, the accuracy remains comparable; for cases where the uninformed FEX converges, both methods yield similar approximation errors, and for the linear conservation law equations, both achieve exact solutions.
-These results highlight the effectiveness of integrating the fine-tuned T5 model into the FEX framework, demonstrating that LLM-informed operator selection leads to a more efficient and reliable symbolic PDE solver.
-3.3.2 Randomly Generated PDEs
-To further validate the reliability of our approach, we evaluated the performance of the FEX algorithm on a set of randomly generated PDE examples.
-For each PDE type, we randomly generated 50 instances using a depth-2 computational tree structure, as described in Section 2.2.
+In this subsection, we evaluate the performance of the fine-tuned BART, T5 and Llama3 models in predicting operator sets for the two PDE types.
+Specifically, we track the average number of mismatched operators on the test dataset for each epoch.
+The computational results are summarized in Figure 3.
+Interested readers are referred to the Table 3 in Appendix C.2 for the detailed computational results.
+Figure 3: Comparison between T5, BART, Llama3-3B and Llama3-8B in terms of average mismatch on test dataset.
+Based on the computational results, LLaMA-8B consistently achieves the lowest number of average mismatch, starting at 0.83 and decreasing to 0.38 within 12 epochs, demonstrating superior performance and training stability.
+The smaller LLaMA-3B variant also performs competitively, with slightly higher values than LLaMA-8B at each epoch, further confirming the effectiveness of the LLaMA architecture.
+These results indicate that, with fine-tuning, LLaMA models can accurately predict the operators in PDE solutions on average.
+In contrast, T5 starts with a significantly higher initial loss of 1.72, converging more slowly to 1.30, suggesting less efficient learning.
+BART begins with the highest initial loss (8.53), which decreases rapidly but still remains substantially higher than that of the LLaMA models by the final epoch (1.65).
+Overall, the decoder-only LLaMA models, particularly LLaMA-8B, exhibit more effective and stable learning dynamics compared to the encoder-decoder models T5 and BART, with larger model sizes empirically yielding better performance.
+4.3 Effectiveness of LLM-informed FEX
+In this section, we compare the performance of LLM-informed FEX with the fine-tuned LLaMA-8B model and the (original) uninformed FEX.
+To this end, we evaluated the performance of the FEX algorithm on a set of 100 randomly generated PDE examples, including 50 Poisson equations and 50 Linear Conservation Law equations.
 Since FEX is a randomized reinforcement learning approach, each instance was solved five times to obtain average computational results, including the number of iterations until convergence, computation time, and solution error.
-The FEX algorithm was limited to a maximum of 150 iterations and an error tolerance of $10^{-3}$ (i.e., $\varepsilon<1\times 10^{-3}$), where $\varepsilon$ was computed during training using a composite loss that combines PDE residuals and boundary violations over a set of sampled points from the domain and its boundary, respectively.
-A PDE instance was considered successfully solved if the desired accuracy was achieved within 150 iterations.
-The computational results are summarized in Tables 6 and 7.
-Table 6: Summary Statistics for Poisson PDEs
-Metric Uninformed FEX LLM-Informed FEX Improvement (Reduction)
-Success Ratio (Valid Solutions) 51.06% (24/47) 68.09% (32/47) $\uparrow 33.35\%$
-For Cases Where BOTH Versions Yield Valid Solutions
-Mean Iterations 61.80 27.87 $\downarrow 54.90\%$
-Mean Duration (min) 40.73 20.25 $\downarrow 50.27\%$
-Table 7: Summary Statistics for Linear Conservation Law PDEs
-Metric Uninformed FEX LLM-Informed FEX Improvement (Reduction)
-Success Ratio (Valid Solutions) 66.67% (30/45) 64.44% (29/45) $\downarrow 3.46\%$
-For Cases Where BOTH Versions Yield Valid Solutions
-Mean Iterations 24.87 13.98 $\downarrow 43.80\%$
-Mean Duration (min) 6.43 3.60 $\downarrow 44.03\%$
-The results in Tables 6 and 7 indicate that the LLM-informed FEX method consistently accelerates convergence compared to the Uninformed FEX, though its impact on solution validity varies across PDE types.
-For Poisson PDEs, the LLM-informed approach significantly improves the success ratio, increasing it by 33.35% (from 51.06% to 68.09%), while also reducing the number of iterations by 54.90% and computation time by 50.27%.
-This suggests that incorporating LLM-informed heuristics enhances both accuracy and efficiency for this class of problems.
-However, for linear conservation law PDEs, the LLM-informed approach slightly reduces the success ratio by 3.46%, indicating that its effectiveness in improving solution validity may be problem-dependent.
-This decline appears to stem from the inherent instability of FEX, even when accurate operators are provided.
-Despite this, the LLM-informed method still achieves a 43.80% reduction in iterations and a 44.03% decrease in computation time, demonstrating substantial gains in computational efficiency.
-Overall, the LLM-informed approach consistently accelerates convergence, making it a valuable tool for solving PDEs more efficiently.
-However, its impact on solution validity remains inconsistent.
-This suggests that while the LLM-informed approach effectively reduce computational costs, one may need to conduct further refinement to the PDE solver itself to improve robustness across different PDE types.
+The detailed computational results can be found in Appendix C.3, and Table 1 and Figure4 summarize the statistics of these results.
+PDE Type Metric Uninformed FEX LLM-Informed FEX Speedup
+Conservation Avg. Iter 28.52 5.23 5.45$\times$
+Avg. Time (m) 23.67 3.95 6.00$\times$
+Poisson Avg. Iter 28.73 6.47 4.44$\times$
+Avg. Time (m) 62.07 14.09 4.41$\times$
+Table 1: Comparison on computational efficiency.
+Figure 4: Comparison on solution accuracy. The accuracy is in log-domain.
+The computational results show that the LLM-informed FEX significantly improves efficiency over the uninformed FEX.
+A key advantage is the reduction in the number of operators used.
+The uninformed FEX always relies on a fixed set of 3 binary and 9 unary operators.
+In contrast, the LLM-informed FEX dynamically and effectively selects smaller and more relevant operator sets.
+This smalle and accurate search space leads to a substantial drop in the number of iterations needed for convergence.
+This efficiency gain further results in significant computational time savings.
+Despite this excellent speedup, the accuracy remains competitive: when the uninformed FEX converges, both methods produce similar approximation errors.
+And the LLM-informed FEX demonstrates better performance in terms of stablility as shown in Figure 4.
+These findings demonstrate that integrating the fine-tuned LLM into symbolic machine learning based PDE solvers enables more efficiency and reliability via informed operator selection.
 5 Conclusions
 In this work, we explored the potential of large language models (LLMs) to uncover symbolic relationships in partial differential equations (PDEs), a largely unexplored challenge in the intersection of AI and scientific computing.
-Our results demonstrate that fine-tuned LLMs can effectively predict the operators involved in PDE solutions by leveraging symbolic information from the governing equations.
-By integrating these predictions into the finite expression method (FEX), we significantly enhanced both the efficiency and accuracy of analytical PDE approximations.
+Based on our theoretical insights on Poisson equation, our results demonstrate that fine-tuned LLMs can effectively predict the operators involved in PDE solutions by leveraging symbolic information from the governing equations.
+By integrating these predictions into symbolic machine learning approaches, such as the finite expression method (FEX), we significantly enhanced both the efficiency and accuracy of analytical PDE approximations.
 Compared to the traditional uninformed FEX, the LLM-informed approach reduces the number of operators required, accelerates convergence, and maintains high solution accuracy, providing a fully interpretable and computationally efficient pipeline.
 These findings highlight the promising role of LLMs in advancing symbolic reasoning for scientific problems, paving the way for further exploration of AI-driven methodologies in mathematical modeling and equation solving.
-Acknowledgement
+Acknowledgments and Disclosure of Funding
 The authors were partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and the DARPA D24AP00325-00.
+Appendix A Binary Expression Trees
+The building blocks of the FEX methods in [30] are illustrated in the following Table 5.
+By increasing the depth of the rexpression tree, the number of functions that can be represented can be significantly increased.
+Also, the mechanism can automatically handle high-dimensional variables.
+Figure 5: Computational expression tree [30].
+Appendix B Proof of Theorem 1
+Let $W_{t},\;t\geq 0$ be a $d$-dimensional Brownian motion such that $W_{t}$ is $F_{t}$-measurable with probability measure $P$. We define the stochastic process
+$X_{t}=X_{0}+W_{t},\quad t\geq 0,$
+and denote by $P_{x}$ the probability measure conditioned on $X_{0}=x$ for any $x\in\mathbb{R}^{d}$, and the expectation with respect to $P_{x}$ is denoted as $\mathbb{E}_{x}[\cdot]$.
+Accordingly, we define the discrete process $\bar{X}_{k},\;k\geq 0$ as follows:
+$\bar{X}_{k}=\bar{X}_{k-1}+Y_{k}\mathcal{D}(\bar{X}_{k-1}),\quad k\geq 1,\quad X_{0}=x,$
+where $Y_{k},\;k\geq 1$ is indpendently, indentically and uniformly distribution on the unit sphere according to $X_{\tau(B(0,1))}$ with respect to a probability distribution $P_{0}$.
+Obviously, $\bar{X}_{k}$ depends on $x,Y_{1},\dots,Y_{k}$, hence to emphasize this dependency, we also dentoe
+$\bar{X}_{k}:=\bar{X}_{k}(x,Y_{1},\dots,Y_{k}),\quad k\geq 1.$
+For any open, nonempty set $\mathcal{X}\subseteq\mathbb{R}^{d}$, we define the first exit time of the process $X_{t}$ starting at $x\in\mathcal{X}$ from $\mathcal{X}$ by
+$\tau(\mathcal{X}):=\inf\left\{t\geq 0\;:\;X_{t}\notin\mathcal{X}\right\}.$
+Then, from [23], we see that for $x\in\Omega$, the solution $u(x)$ for the Poisson equation admits the following expression:
+[…]
+We first consider approximating the first term in (5) using the Monte-Carlo sampling technique. To this end, we define the function $g^{(M,\{K_{i}\})}:\mathbb{R}^{d}\to\mathbb{R}$ as
+$g^{(M,\{K_{i}\})}(x):=\frac{1}{M}\sum_{i=1}^{M}g\left(\bar{X}_{K_{i}}(x,Y_{i,1},\dots,Y_{i,K_{i}})\right),$
+where $M\geq 1$, $K_{i}\geq 1$ for $i=1,\dots,M$, and $Y_{i,k},\;k=1,\dots,K_{i}$ are unit vectors.
+Then, for any $\delta\in(0,1)$, there exist $M$ and $K_{i}$, $1\leq i\leq M$, all polynomial in $\delta^{-1}$ [15], such that
+[…]
+Similarly, we consider approximating the second term in (5) using the Monte-Carlo sampling technique.
+Let $Y_{i,k},\;k=1,\dots,K_{i}$ and $y_{i,j,k},\;k=1,\dots,K_{i},\;j=1,\dots,J$ be random unit vectors, for $i=1,\dots,M$, with $M,K_{i},J$ being positive integers, then for any $\delta>0$, one can show that the following function $f^{(M,\{K_{i}\},J)}:\mathbb{R}^{d}\to\mathbb{R}$ defined by
+$f^{(M,\{K_{i}\},J)}(x):=$ $\;\frac{1}{MJ}\sum_{i=1}^{M}\sum_{k=1}^{M_{i}}\sum_{j=1}^{J}\mathcal{D}\left(\bar{X}_{k}(x,Y_{i,1},\dots,Y_{i,k})\right)^{2}$
+×$\qquad\times f\left(\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1})+\mathcal{D}(\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1}))y_{i,j,k}\right),$
+can approximate the second term (5) in sense that
+[…]
+if one chooses suitable $M,\;J$ and $K_{i},\;i=1,\dots,M$ all polynomial in $\delta^{-1}$ [15]. Therefore, the proof is completed.
+Appendix C Detailed Computational Results
+C.1 Illustrative Example for Binary Encoding
+Consider the following two functions:
+$h_{1}(x_{1},x_{2}):=(5x_{1})^{2}+\sin(3x_{2})\cdot x_{2},\quad h_{2}(x_{1}):=5\exp(2x_{1})\cdot(\cos(6x_{1}))^{3}.$
+The binary representation of them is provided in Table 2.
+Binary Vector
+Function Preprocessed Operator Set
+$x_{1}$
+$x_{2}$
+^2
+^3
+SIN
+COS
+EXP
+$h_{1}(x_{1},x_{2})$ [ $x_{1}$, $x_{2}$, ^2, +, *, SIN ] 1 1 1 0 1 1 1 0 0
+$h_{2}(x_{1})$ [ $x_{1}$, ^3, *, COS, EXP ] 1 0 0 1 0 1 0 1 1
+Table 2: Sample binary-vector encoding for two example expressions. The size of the dictionary is 9.
+The resulted mismatch can then be computed as $7$, indicating that the two functions are very different from each other.
+C.2 Detailed Computational Results for Average Mismatch
+Epoch LLaMA-8B LLaMA-3B T5 BART
+1 0.82697 0.80510 1.71635 8.52542
+2 0.65606 0.65040 1.40708 5.61399
+3 0.55470 0.59328 1.37730 4.69332
+4 0.51061 0.53747 1.37543 2.26791
+5 0.48742 0.48601 1.32494 1.72084
+6 0.45515 0.47934 1.36476 1.88165
+7 0.42818 0.45894 1.34016 1.58241
+8 0.41510 0.44808 1.32464 2.07886
+9 0.39616 0.42293 1.31991 1.65313
+10 0.39470 0.41894 1.30461 1.48738
+11 0.38308 0.40672 1.30633 1.74097
+12 0.38394 0.40697 1.29627 1.64541
+Table 3: Comparison of Average Mismatch Across Models During Fine-tuning on 198K Dataset
+C.3 Detailed computational results for FEX
+True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $2\bigl{(}-2x_{2}^{3}+2\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\exp\bigl{(}4x_{2}\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8x_{2}^{4}-2\bigl{(}2\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4x_{2}^{4}+4x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\exp(x_{1})+2\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\sin\bigl{(}2x_{3}^{3}-2\exp(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4x_{2}^{4}+4x_{3}^{4}+8\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\sin\bigl{(}2x_{3}^{3}+2\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\exp\bigl{(}2x_{2}^{2}-2\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\sin^{2}(x_{2})\cos^{2}(x_{1})$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\exp\bigl{(}2\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\sin^{4}(x_{1})\cos^{4}(x_{1})$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\exp(2x_{2})-2\exp\bigl{(}2x_{1}^{2}-2x_{2}^{2}\bigr{)}$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\exp\bigl{(}2x_{3}^{4}-2\exp(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2x_{2}^{4}+2x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2x_{2}^{4}-2x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\sin\bigl{(}2\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\exp\bigl{(}2\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\cos\bigl{(}4x_{2}^{4}\bigr{)}\cos\bigl{(}2x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\sin\bigl{(}4\exp(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2x_{3}^{4}-2\bigr{)}^{3}+2\cos\bigl{(}2x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{4}+2\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $22\sin\bigl{(}2x_{1}^{4}-2\sin(x_{3})\bigr{)}-$ 2 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $22\sin\bigl{(}2x_{3}+2\bigr{)}-$ 2 2.2 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\cos\bigl{(}2\sin(x_{2})+2\cos(x_{3})\bigr{)}$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\cos\bigl{(}2x_{1}^{3}-2x_{3}^{2}\bigr{)}$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{2}+2\sin(x_{3})\bigr{)}+$ 2 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $232x_{3}^{6}+$ 2 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$ $2\bigl{(}-2x_{2}^{3}+2\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\exp\bigl{(}4x_{2}\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8x_{2}^{4}-2\bigl{(}2\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4x_{2}^{4}+4x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\exp(x_{1})+2\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\sin\bigl{(}2x_{3}^{3}-2\exp(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4x_{2}^{4}+4x_{3}^{4}+8\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\sin\bigl{(}2x_{3}^{3}+2\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\exp\bigl{(}2x_{2}^{2}-2\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\sin^{2}(x_{2})\cos^{2}(x_{1})$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\exp\bigl{(}2\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\sin^{4}(x_{1})\cos^{4}(x_{1})$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\exp(2x_{2})-2\exp\bigl{(}2x_{1}^{2}-2x_{2}^{2}\bigr{)}$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\exp\bigl{(}2x_{3}^{4}-2\exp(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2x_{2}^{4}+2x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2x_{2}^{4}-2x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\sin\bigl{(}2\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\exp\bigl{(}2\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\cos\bigl{(}4x_{2}^{4}\bigr{)}\cos\bigl{(}2x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\sin\bigl{(}4\exp(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2x_{3}^{4}-2\bigr{)}^{3}+2\cos\bigl{(}2x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{4}+2\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$
+Table 4: Results of LLM-informed and original FEX runs on 50 randomly generated Linear Conservation Law functions.
+True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $8\,x_{1}\,\mathrm{exp}(x_{2})$ 42.4 1.2 74.44 0.5900 $3.75\times 10^{-4}$ $4.91\times 10^{-4}$ $2128\,x_{1}^{12}+$ 2 40.2 1.2 70.704 0.516 $3.97\times 10^{-2}$ $2.14\times 10^{-2}$ $-32\,x_{2}^{10}$ 35.2 1.2 71.586 0.536 $7.44\times 10^{-3}$ $6.71\times 10^{-3}$ $-8\,x_{1}^{4}\,x_{2}$ 25.0 1.2 53.342 0.510 $9.18\times 10^{-4}$ $3.86\times 10^{-4}$ $2\,\cos\bigl{(}4\,x_{2}^{6}\bigr{)}$ 44.2 3.0 85.792 4.326 $1.68\times 10^{-2}$ $1.23\times 10^{-2}$ $-8\,x_{2}\,\cos(x_{2})+2$ 32.4 2.4 78.076 4.548 $3.78\times 10^{-4}$ $3.06\times 10^{-4}$ $28\,x_{1}^{2}\,\cos(x_{2})+$ 2 42.8 3.2 85.55 6.154 $2.05\times 10^{-4}$ $1.85\times 10^{-4}$ $-4\,\mathrm{exp}(x_{3})+4\,\sin(x_{1})+2$ 20.2 1.6 43.712 2.150 $1.14\times 10^{-4}$ $2.04\times 10^{-4}$ 2($22\bigl{(}2\,x_{2}^{2}-2\,x_{3}^{4}\bigr{)}^{2}+$ 2 31.8 3.0 69.054 3.400 $6.55\times 10^{-3}$ $2.66\times 10^{-3}$ $-32\,\sin(x_{2})^{4}$ 15.6 1.6 36.210 1.704 $4.34\times 10^{-3}$ $4.66\times 10^{-3}$ $-4\,x_{2}^{3}-4\,x_{3}^{4}+16$ 20.6 2.4 45.110 2.716 $3.88\times 10^{-4}$ $4.68\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}-2\,\mathrm{exp}(x_{3})+2\,\cos(x_{3})\bigr{)}+$ 2 22.0 2.8 52.848 4.994 $3.83\times 10^{-4}$ $4.28\times 10^{-4}$ $28\,x_{1}^{2}\,x_{3}+$ 2 11.4 1.6 26.022 1.742 $3.58\times 10^{-4}$ $1.82\times 10^{-4}$ $2\,\sin\bigl{(}2\,\sin(x_{1})-2\,\cos(x_{3})\bigr{)}$ 19.2 3.0 47.168 4.768 $7.78\times 10^{-5}$ $3.13\times 10^{-5}$ $4\bigl{(}2\,\cos(x_{2})-2\,\cos(x_{3})\bigr{)}^{4}$ 10.2 1.6 21.852 1.476 $3.80\times 10^{-3}$ $3.12\times 10^{-3}$ 2($22\bigl{(}2\,x_{1}+2\,\mathrm{exp}(x_{1})\bigr{)}^{2}-$ 2 26.4 4.2 44.664 6.108 $1.07\times 10^{-3}$ $2.94\times 10^{-3}$ $22\,\sin\bigl{(}2\,x_{2}^{3}+2\,x_{3}^{4}\bigr{)}+$ 2 29.0 4.8 73.818 12.376 $6.12\times 10^{-4}$ $3.11\times 10^{-4}$ 2($22\bigl{(}-2\,x_{1}+2\,\mathrm{exp}(x_{3})\bigr{)}^{2}-$ 2 20.4 3.4 35.796 4.762 $1.96\times 10^{-3}$ $2.52\times 10^{-3}$ $22\,\cos\bigl{(}2\,\mathrm{exp}(x_{1})+2\,\sin(x_{2})\bigr{)}+$ 2 69.6 11.6 157.764 27.654 $3.92\times 10^{-5}$ $1.17\times 10^{-4}$ $2\,\sin\bigl{(}2\,\sin(x_{2})-2\,\cos(x_{2})\bigr{)}$ 18.2 3.2 37.792 5.254 $1.01\times 10^{-5}$ $3.35\times 10^{-5}$ $2\bigl{(}-2\,x_{2}^{3}+2\,\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\,\mathrm{exp}\bigl{(}4\,x_{2}\,\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8\,x_{2}^{4}-2\bigl{(}2\,\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4\,x_{2}^{4}+4\,x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\,\mathrm{exp}(x_{1})+2\,\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\,\sin\bigl{(}2\,x_{3}^{3}-2\,\mathrm{exp}(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4\,x_{2}^{4}+4\,x_{3}^{4}+8\,\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\,\sin\bigl{(}2\,x_{3}^{3}+2\,\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\,\mathrm{exp}\bigl{(}2\,x_{2}^{2}-2\,\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\,\sin(x_{2})^{2}\,\cos(x_{1})^{2}$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\,\mathrm{exp}\bigl{(}2\,\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\,\sin(x_{1})^{4}\,\cos(x_{1})^{4}$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\,\mathrm{exp}(2\,x_{2})-2\,\mathrm{exp}(2\,x_{1}^{2}-2\,x_{2}^{2})$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\,\mathrm{exp}\bigl{(}2\,x_{3}^{4}-2\,\mathrm{exp}(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2\,x_{2}^{4}+2\,x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2\,x_{2}^{4}-2\,x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2\,x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\,\sin\bigl{(}2\,\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\,\mathrm{exp}\bigl{(}2\,\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\,\cos\bigl{(}4\,x_{2}^{4}\bigr{)}\,\cos\bigl{(}2\,x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\,\sin\bigl{(}4\,\mathrm{exp}(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2\,x_{3}^{4}-2\bigr{)}^{3}+2\,\cos\bigl{(}2\,x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\,\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}2\,x_{1}^{4}+2\,\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $22\,\sin\bigl{(}2\,x_{1}^{4}-2\,\sin(x_{3})\bigr{)}-$ 2 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $22\,\sin\bigl{(}2\,x_{3}+2\bigr{)}-$ 2 2.2 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\,\cos\bigl{(}2\,\sin(x_{2})+2\,\cos(x_{3})\bigr{)}$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\,\cos\bigl{(}2\,x_{1}^{3}-2\,x_{3}^{2}\bigr{)}$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}2\,x_{1}^{2}+2\,\sin(x_{3})\bigr{)}+$ 2 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $232\,x_{3}^{6}+$ 2 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$
+Table 5: Results of LLM-informed and original FEX runs on 50 randomly generated Poisson functions.

### v3

@@
-From Equations to Insights: Unraveling Symbolic Structures in PDEs with LLMs
+
 Abstract
@@
 This work opens new avenues for understanding the symbolic structure of scientific problems and advancing their solution processes.
+keywords:
+Large Language Models, Finite Expression Method, High Dimensional PDEs, Operator Relation
+{MSCcodes}
+65N75, 68T20, 90C15
 1 Introduction
-Numerous grand challenges in scientific fields, including physics, engineering, biology, and chemistry, can be modeled as partial differential equations (PDEs), which relate unknown functions defined on a multidimensional domain (e.g., in terms of space and time) to its rates of change with respect to each of those variables [11].
-Mathematically, a PDE takes the following form
-$Du=f,\quad\text{in }\Omega,\quad Bu=g,\quad\text{on }\partial\Omega,$ (1)
-where $D$ denotes a differential operator defining the dynamics, $Bu=g$ represents the boundary condition (e.g., Dirichlet, Neumann, and Robin), and $\Omega\subseteq\mathbb{R}^{d}$ is the domain with boundary $\partial\Omega$.
+Partial differential equations (PDEs) are the mathematical language of the natural sciences: they encode how fields evolve and interact by linking an unknown function on space–time to its derivatives.
+From turbulent flows and elastic structures to reaction–diffusion in biology and quantum systems, many grand scientific challenges are governed by PDE models [12].
+Mathematically, a strong-form PDE constrains an unknown function $u:\Omega\to\mathbb{R}$ on a domain $\Omega\subseteq\mathbb{R}^{d}$ by relating $u$ to its partial derivatives together with boundary data
+(1) $Du=f,\quad\text{in }\Omega,\quad Bu=g,\quad\text{on }\partial\Omega.$
+Here $D$ is a differential operator of order $k$ that may depend on $x$, $u$, and derivatives of $u$ up to order $k$; a linear prototype is $Du=\sum_{|\alpha|\leq k}a_{\alpha}(x)\,D^{\alpha}u$ with multi-index $\alpha=(\alpha_{1},\ldots,\alpha_{d})$ and $D^{\alpha}u:=\partial^{|\alpha|}u/\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}$, while $f:\Omega\to\mathbb{R}$ is a given source term and $g:\partial\Omega\to\mathbb{R}$ is given.
+The boundary operator $B$ prescribes data on $\partial\Omega$.
+Some commonly used boundary conditions include $Bu=\gamma_{0}u=u|_{\partial\Omega}=g$ (Dirichlet), $Bu=\gamma_{1}u=\partial_{\boldsymbol{n}}u=g$ with $\boldsymbol{n}$ the outward unit normal (Neumann), or $Bu=\gamma_{1}u+\kappa\,\gamma_{0}u=g$ for $\kappa\geq 0$ (Robin) where $\gamma_{0}:H^{1}(\Omega)\to H^{1/2}(\partial\Omega)$ and $\gamma_{1}:H^{1}(\Omega)\to H^{-1/2}(\partial\Omega)$ denote continuous trace maps.
+For time-dependent problems on $Q=(0,T)\times\Omega$, one augments the system with the initial condition $u(0,\cdot)=u_{0}$ in $\Omega$.
+Classical examples include the parabolic and hyperbolic models
+$u_{t}-\nabla\!\cdot(A\nabla u)=f,\qquad u_{tt}-\nabla\!\cdot(A\nabla u)=f,$
+where $A$ denotes the diffusion/conductivity tensor.
+When classical derivatives are unavailable, the statement $Du=f$ and $Bu=g$ is interpreted in a weak (variational) sense on Sobolev spaces, where existence and uniqueness follow under standard boundedness and coercivity assumptions (e.g., by the Lax–Milgram theorem).
+And the principal symbol of $D$ then classifies the PDE as elliptic, parabolic, or hyperbolic.
 PDEs often arise in high-dimensional settings (i.e., $d$ is large); for example, well-known instances such as the Poisson equation [47], the linear conservation law [6], and the nonlinear Schrödinger equation [18] naturally involve multiple spatial and temporal variables.
@@
 However, the underlying symbolic relationship among $u$, $f$ and $g$ often remains underexplored.
-To address this gap, we pose the following question:
-Assuming that the PDE (1) admits an analytical solution $u$, and given the operators in $f$ and $g$, how are the operators in $u$ related to those in $f$ and $g$?
+To address this gap, we pose the following question: Assuming that the PDE (1) admits an analytical solution $u$, and given the operators in $f$ and $g$, how are the operators in $u$ related to those in $f$ and $g$?
 Understanding this symbolic connection is crucial for deriving fully interpretable, closed-form solutions, a feature lacking in most existing solution methods, particularly in the high-dimensional setting.
 As an illustrative example, we consider solving the PDE (1) via the least square method [10, 25, 41], which defines a straightforward loss to characterize the error of the estimated solution by
-$L(u)=\|Du-f\|^{2}_{L^{2}(\Omega)}+\lambda\|Bu-g\|^{2}_{L^{2}(\partial\Omega)},$ (2)
+(2) $L(u)=\|Du-f\|^{2}_{L^{2}(\Omega)}+\lambda\|Bu-g\|^{2}_{L^{2}(\partial\Omega)},$
 where $\lambda>0$ balances the influence of boundary conditions. The goal is to find a function $u^{*}\in L^{2}(\Omega)$ so that the least square loss is minimized, i.e.,
-$u^{*}=\mathrm{argmin}\left\{L(u)\;:\;u\in L^{2}(\Omega)\right\}.$ (3)
+(3) $u^{*}=\mathrm{argmin}\left\{L(u)\;:\;u\in L^{2}(\Omega)\right\}.$
 Because the search space $L^{2}(\Omega)$ is an infinite dimensional space, finding the solution $u^{*}$ is extremely challenging.
@@
 The process begins with the generation of a comprehensive, structured dataset composed of symbolic expressions derived from a diverse range of PDE types, including elliptic, parabolic, and hyperbolic equations, along with their associated boundary and initial conditions.
-These symbolic expressions are represented as computational trees (see Figure 5), where nodes are populated with unary (e.g., sin, cos, exp) and binary operators (e.g., +, -, *, /).
+These symbolic expressions are represented as computational trees (see Fig.
+1), where nodes are populated with unary (e.g., sin, cos, exp) and binary operators (e.g., +, -, *, /).
 This tree structure captures the hierarchical nature of mathematical expressions and ensures a systematic representation of the PDE solutions.
+Figure 1: Computational expression tree [32].
 To prepare the data for LLM processing, the symbolic expressions are tokenized and converted into postfix notation (Reverse Polish Notation).
@@
 More importantly, it reflects the relative positional relationship of operators in the expression.
-The tokenized dataset is then used to fine-tune foundational LLMs such as T5, BART, and Llama.
+The tokenized dataset is then used to fine-tune foundational LLMs.
 During fine-tuning, the models are trained to predict the sequence of operators that constitute the symbolic solution to a new PDE.
@@
 The predicted operator sets are highly refined, containing only the operators necessary to construct the solution, thereby eliminating redundant or irrelevant operators.
-This refinement is crucial for the subsequent application of the finite-expression method (FEX), a reinforcement learning based symbolic regression technique that searches for the optimal combination of operators and constants to approximate the PDE solution.
+This refinement is crucial for the subsequent application of the finite-expression method (FEX), a reinforcement learning–based symbolic regression technique that searches for the optimal combination of operators and constants to approximate the PDE solution.
 By providing FEX with a reduced and targeted operator set, the combinatorial search space is significantly narrowed.
 This reduction not only enhances the accuracy of the solution by focusing on relevant operators, but also dramatically improves computational efficiency by reducing the number of potential combinations that need to be evaluated.
+While we adopt FEX in this work, it is not the only viable approach.
+If the symbolic relationship between a PDE and its solution can be effectively learned, then a wide range of symbolic regression methods [26], such as evolutionary algorithms [38], can be directly applied using the learned operator set.
 In summary, the integration of LLMs into this workflow transforms the process of symbolic PDE solving.
-By generating structured datasets, fine-tuning models to predict operator sequences, and leveraging these predictions to guide FEX, we achieve a more accurate and computationally efficient approach to discovering symbolic solutions for complex PDEs.
-This methodology bridges the gap between machine learning and symbolic mathematics, offering great potential for scientific computing and engineering applications.
+By generating structured datasets, fine-tuning models to predict operator sequences, and leveraging these predictions to guide PDE solvers, we achieve a more accurate and computationally efficient approach to discovering symbolic solutions for complex PDEs.
+This methodology bridges the gap between machine learning and symbolic mathematics, offering great potential for scientific and engineering applications.
 Particularly, our contributions are summarized as follows:
-(1) We formally pose the problem of discovering the symbolic relationship between the solution of a PDE and the associated problem data, including the boundary conditions and right-hand-side functions.
+  1. 1.
+    PDE symbolic relations.
+We formally pose the problem of discovering the symbolic relationship between the solution of a PDE and the associated problem data, including the boundary conditions and source functions.
 Establishing this connection can provide deeper insights into the underlying structure of PDEs, enabling more precise analytical solutions and facilitating the development of efficient numerical methods which provides high-quality analytical solutions.
-By highlighting the potential impact of such relationships, our work lays the foundation for further studies in this area, aiming to bridge the gap between problem formulation and solution representation.
-(2) We propose fine-tuning LLMs to learn the symbolic relationship between problem data and the solution of a PDE, demonstrating remarkable effectiveness in predicting the operators that characterize the solution.
-By integrating this approach with the FEX method, our numerical findings show that while LLMs may not be inherently suited for directly solving PDEs or other complex scientific problems, they provide substantial benefit in identifying underlying structures and symbolic patterns.
-This capability significantly streamlines and enhances the solution process.
-Our work underscores the potential of LLMs as transformative tools in the analytical and computational study of PDEs, bridging the gap between symbolic reasoning and sophisticated mathematical problem-solving.
+By highlighting the potential impact of such relationships, our work lays the foundation for further studies, aiming to bridge the gap between problem formulation and solution representation.
+  2. 2.
+    LLM-guided symbolic PDE solvers.
+We fine-tune LLMs to predict the operator set that captures the solution’s structure and couple these predictions with the FEX pipeline to search over symbolic programs that produce fully interpretable surrogates $u_{\theta}$.
+Experiments show that, while LLMs are not direct PDE solvers, their operator- and geometry-aware priors sharply prune the search over expressions, improving FEX’s sample efficiency and stability and yielding higher-quality closed-form or near–closed-form solutions across PDE families.
+The same operator-selection front end is modular and can be integrated with other symbolic PDE solvers where operator choice is the critical bottleneck.
+  3. 3.
+    Theory for policy-gradient–based symbolic learning.
+We provide convergence guarantees for a stochastic projected policy-gradient method that maximizes the physics-informed return over symbolic programs.
+In particular, under standard conditions, we prove that projected updates converge to first-order stationary policies and establish a sample complexity of $\mathcal{O}(\epsilon^{-4})$ to reach an $\epsilon$-stationary point in expectation, thereby placing the LLM+FEX pipeline on firm theoretical footing among state-of-the-art alternatives.
 1.1 Related work
@@
 Specifically, we focus on the classical Poisson equation with Dirichlet boundary condition
-$-\Delta u=f\quad\text{in }\Omega,\qquad u=g\quad\text{on }\partial\Omega,$ (4)
+(4) $\Delta u=f\quad\text{in }\Omega,\qquad u=g\quad\text{on }\partial\Omega,$
 where $\Omega\subseteq\mathbb{R}^{d}$ is a compact domain.
 By analyzing this simple class of elliptic PDEs under, we aim to shed light on the structural properties and solution behavior relevant to the broader question, thereby laying the groundwork for more comprehensive theoretical development in future work.
-Theorem 1.
+Theorem 2.1.
 Let $\Omega\subset\mathbb{R}^{d}$ be a compact domain with boundary $\partial\Omega$ such that the distance function $\mathcal{D}(x):=\mathrm{dist}(x,\partial\Omega)$ admits an explicit analytical expression.
@@
   2. 2.
-    $\tilde{u}$ is constructed using only the elementary arithmetic operators +,×$+,\;\times,\;(\cdot)^{2}$ and the operators appearing in the evaluation of $\mathcal{D}(\cdot)$, $f(\cdot)$, and $g(\cdot)$,
+    $\tilde{u}$ is constructed using only the elementary arithmetic operators $\times,\;(\cdot)^{2}$ and the operators appearing in the evaluation of $\mathcal{D}(\cdot)$, $f(\cdot)$, and $g(\cdot)$,
   3. 3.
@@
     where $\texttt{poly}(\delta^{-1})$ denotes a quantity polynomial in $\delta^{-1}$.
-Proof.
-See Appendix B. ∎
+Proof 2.2.
+Let $W_{t},\;t\geq 0$ be a $d$-dimensional Brownian motion such that $W_{t}$ is $F_{t}$-measurable with probability measure $P$. We define the stochastic process
+$X_{t}=X_{0}+W_{t},\quad t\geq 0,$
+and denote by $P_{x}$ the probability measure conditioned on $X_{0}=x$ for any $x\in\mathbb{R}^{d}$, and the expectation with respect to $P_{x}$ is denoted as $\mathbb{E}_{x}[\cdot]$.
+Accordingly, we define the discrete process $\bar{X}_{k},\;k\geq 0$ as follows:
+$\bar{X}_{k}=\bar{X}_{k-1}+Y_{k}\mathcal{D}(\bar{X}_{k-1}),\quad k\geq 1,\quad X_{0}=x,$
+where $Y_{k},\;k\geq 1$ is indpendently, indentically and uniformly distribution on the unit sphere according to $X_{\tau(B(0,1))}$ with respect to a probability distribution $P_{0}$.
+Obviously, $\bar{X}_{k}$ depends on $x,Y_{1},\dots,Y_{k}$, hence to emphasize this dependency, we also dentoe
+$\bar{X}_{k}:=\bar{X}_{k}(x,Y_{1},\dots,Y_{k}),\quad k\geq 1.$
+For any open, nonempty set $\mathcal{X}\subseteq\mathbb{R}^{d}$, we define the first exit time of the process $X_{t}$ starting at $x\in\mathcal{X}$ from $\mathcal{X}$ by
+$\tau(\mathcal{X}):=\inf\left\{t\geq 0\;:\;X_{t}\notin\mathcal{X}\right\}.$
+Then, from [24], we see that for $x\in\Omega$, the solution $u(x)$ for the Poisson equation admits the following expression:
+(5) $\;u(x)=\mathbb{E}_{x}\left[g(X_{\tau(\Omega)})\right]+\frac{1}{2}\mathbb{E}_{x}\left[\sum_{k\geq 1}\mathcal{D}(\bar{X}_{k})^{2}E_{0}\right],$
+where
+$E_{0}:=\mathbb{E}_{0}\left[\int_{0}^{\tau(B(0,1))}f\left(\bar{X}_{k-1}+D(\bar{X}_{k-1})\cdot t\right)dt\right].$
+We first consider approximating the first term in (5) using the Monte-Carlo sampling technique. To this end, we define the function $g^{(M,\{K_{i}\})}:\mathbb{R}^{d}\to\mathbb{R}$ as
+$g^{(M,\{K_{i}\})}(x):=\frac{1}{M}\sum_{i=1}^{M}g\left(\bar{X}_{K_{i}}(x,Y_{i,1},\dots,Y_{i,K_{i}})\right),$
+where $M\geq 1$, $K_{i}\geq 1$ for $i=1,\dots,M$, and $Y_{i,k},\;k=1,\dots,K_{i}$ are unit vectors.
+Then, for any $\delta\in(0,1)$, there exist $M$ and $K_{i}$, $1\leq i\leq M$, all polynomial in $\delta^{-1}$ [16], such that
+$\sqrt{\int_{\Omega}\left(\mathbb{E}_{x}\left[g(X_{\tau(\Omega)})\right]-g^{(M,\{K_{i}\})}(x)\right)^{2}dx}\leq\delta.$
+Similarly, we consider approximating the second term in (5) using the Monte-Carlo sampling technique.
+Let $Y_{i,k},\;k=1,\dots,K_{i}$ and $y_{i,j,k},\;k=1,\dots,K_{i},\;j=1,\dots,J$ be random unit vectors, for $i=1,\dots,M$, with $M,K_{i},J$ being positive integers, then for any $\delta>0$, one can show that the following function $f^{(M,\{K_{i}\},J)}:\mathbb{R}^{d}\to\mathbb{R}$ defined by
+$f^{(M,\{K_{i}\},J)}(x)=\frac{1}{MJ}\sum_{i=1}^{M}\sum_{k=1}^{M_{i}}\sum_{j=1}^{J}\mathcal{D}\left(\bar{X}_{k}(x,Y_{i,1},\dots,Y_{i,k})\right)^{2}\times f\left(Z_{k-1}\right),$
+where
+$Z_{k-1}=\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1})+\mathcal{D}(\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1}))y_{i,j,k},$
+can approximate the second term (5) in sense that
+$\sqrt{\int_{\Omega}\left(\mathbb{E}_{x}\left[\sum_{k\geq 1}\mathcal{D}(\bar{X}_{k})^{2}E_{0}\right]-f^{(M,\{K_{i}\},J)}(x)\right)^{2}dx}\leq\delta,$
+where
+$E_{0}:=\mathbb{E}_{0}\left[\int_{0}^{\tau(B(0,1))}f\left(\bar{X}_{k-1}+\mathcal{D}(\bar{X}_{k-1})\cdot t\right)dt\right],$
+if one chooses suitable M,JM,\;J and $K_{i},\;i=1,\dots,M$ all polynomial in $\delta^{-1}$ [16]. Therefore, the proof is completed.
+The objective of the above theorem is to establish a proof-of-concept framework to guide symbolic discovery of PDE solutions using modern machine learning approaches, including LLMs.
 This theorem offers a partial constructive resolution to the open research question posed in the previous section: it identifies the specific operators needed to represent a $\delta$-accurate approximation of the PDE solution $u$ in terms of the operators used to define the source term $f$, the boundary condition $g$, and the domain geometry via the distance function $\mathcal{D}$.
@@
 Taken together, these insights lay a foundation for a broader theory of operator-preserving approximations for general PDEs, with significant potential impact in scientific machine learning, computational physics, and algorithmic model design.
+We plan to pursue a unified analysis for more general classes of PDEs in future work, and our current numerical results on linear conservation laws already provide clear evidence supporting this direction.
 3 Methodology
@@
 3.1 Binary Computational Trees for Expressions
-Expressions in PDEs are represented as binary computational trees, as depicted in Figure 5 in Appendix A.
+Expressions in PDEs are represented as binary computational trees, as depicted in Figure 1.
 These trees provide a structured and hierarchical way to encode mathematical expressions, enabling efficient exploration and manipulation of the solution space.
@@
 Each node in the computational tree encodes an operator, which may be either a unary or binary operator.
-Unary Operators: Unary operators act on a single operand and include a variety of nonlinear transformations such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other elementary operations (e.g., $\sqrt{\cdot}$, ||$\cdot$|| ⋅ |, and power functions $(\cdot)^{k}$ for $k\in\mathbb{Z}$).
+Unary Operators: Unary operators act on a single operand and include a variety of nonlinear transformations such as trigonometric functions ($\sin(\cdot)$, $\cos(\cdot)$, $\tan(\cdot)$), exponential functions ($\exp(\cdot)$), logarithmic functions ($\lg(\cdot)$, $\ln(\cdot)$), and other elementary operations (e.g., $\sqrt{\cdot}$, $|\cdot|$, and power functions $(\cdot)^{k}$ for $k\in\mathbb{Z}$).
 Each unary operator node is augmented with two scalar parameters, $\alpha$ and $\beta$, which enable the composition of the operator with an affine transformation, yielding expressions of the form $\alpha\cdot u(\cdot)+\beta$, where $u(\cdot)$ denotes the unary operation.
 This parameterization allows the tree to flexibly represent both nonlinear and linearly adjusted transformations, thereby enhancing its expressiveness.
-Binary Operators: Binary operators operate on two operands and include the standard arithmetic operations: addition ($+$), subtraction (--), multiplication ($\times$), and division (///).
+Binary Operators: Binary operators operate on two operands and include the standard arithmetic operations: addition (++), subtraction (-), multiplication ($\times$), and division (//).
 Each binary operator node combines the outputs of its left and right child subtrees to produce more complex expressions.
@@
 3.3 Fine-tuning LLM
-With a generated data set, we can then fine-tune a certain LLM to learn the symbolic relastion in PDEs.
+With a generated data set, we can then fine-tune a certain LLM to learn the symbolic relation in PDEs.
 The whole pipeline is summarized in Figure 3, and we will describe each component in detail for the rest of this subsection.
@@
 Specifically, each data point (in raw text) is serialized into a prompt followed by its target operator sequence, as demonstrated in the following.
-$\underbrace{\texttt{Type:}<\mathrm{PDE}>\,\texttt{|}\,\texttt{RHS:}<\mathrm{ops}>\,\texttt{|}\,\texttt{BC\_Type:}<\mathrm{bc}>\,\texttt{|}\,\texttt{Solution:}}_{\text{PROMPT}}\;\|\|\underbrace{\texttt{<op}_{1}\ \dots\ \texttt{op}_{n}\,\texttt{<EOS>}\texttt{>}}_{\text{TARGET}}$
+$\;\underbrace{\texttt{Type:}<\mathrm{PDE}>\,\texttt{|}\,\texttt{RHS:}<\mathrm{ops}>\,\texttt{|}\,\texttt{BC\_Type:}<\mathrm{bc}>\,\texttt{|}\,\texttt{Solution:}}_{\text{PROMPT}}$
+$\;\underbrace{\texttt{<op}_{1}\ \dots\ \texttt{op}_{n}\,\texttt{<EOS>}\texttt{>}}_{\text{TARGET}}$
+In this work, we intentionally adopted a simple and minimal prompt design to isolate the core capabilities of the fine-tuned language model and avoid introducing prompt-specific biases.
+However, incorporating explicit task descriptions or more structured prompts could potentially enhance performance, and exploring such prompt engineering strategies is a valuable research direction.
 Next, a pretrained tokenizer is applied to convert the input text into its corresponding token sequence.
@@
 The decoder will auto-regressively generate a sequence of tokens that represents the solution in postfix notation, encoding the sequence of mathematical operators and operands that define the PDE solution.
-Since the model’s raw output is a token sequence, a post-processing pipeline is required to extract a clean and interpretable set of predicted operators.
-The first step involves identifying the unique set of operators by removing duplicate tokens from the generated sequence.
+Since the model’s raw output is a token sequence, a post-processing pipeline then extracts a clean and interpretable set of predicted operators.
+The first step involves identifying the unique set of operators by removing duplicate tokens from the sequence.
 Following this, we perform error checking to detect and eliminate misspelled or malformed tokens that may result from tokenization artifacts or character-level prediction errors (issues that can arise because the model learns statistical patterns rather than enforcing strict syntactic correctness).
@@
 Any invalid or unrecognized tokens are subsequently discarded to ensure the integrity of the final operator set.
-To quantitatively evaluate the model’s predictive accuracy, we compute the squared $\ell_{2}$ norm of the difference between the predicted and ground-truth operator sets.
+To measure the model’s predictive accuracy, we compute the squared $\ell_{2}$-norm of the difference between the predicted and ground-truth operator sets.
 Specifically, we encode each operator set as a binary vector over a fixed dictionary of $n$ possible operators, where each vector component indicates the presence or absence of a given operator.
 This encoding transforms the comparison into a well-defined vector space problem.
-An illustrative example of this binary encoding scheme can be found in Appendix C.1.
+An illustrative example of this binary encoding scheme with
+$h_{1}(x_{1},x_{2}):=(5x_{1})^{2}+\sin(3x_{2})\cdot x_{2},\quad h_{2}(x_{1}):=5\exp(2x_{1})\cdot(\cos(6x_{1}))^{3},$
+is showed in Table 1.
 Then, the squared distance between two such operator sets, represented by two binary vectors $y\in\mathbb{R}^{n}$ and $z\in\mathbb{R}^{n}$, can be defined as $\|y-z\|^{2}=\sum_{i=1}^{n}\left(y_{i}-z_{i}\right)^{2}.$ Clearly, it measures the number of mismatched operators between the two operator sets.
+The resulted mismatch can then be computed as $7$, indicating that the two functions are very different from each other.
+Binary Vector
+Operator Set
+$x_{1}$
+$x_{2}$
+ˆ2
+ˆ3
+SIN
+COS
+EXP
+$h_{1}$ [ $x_{1}$, $x_{2}$, ˆ2, +, *, SIN ] 1 1 1 0 1 1 1 0 0
+$h_{2}$ [ $x_{1}$, ˆ3, *, COS, EXP ] 1 0 0 1 0 1 0 1 1
+Table 1: Sample binary‐vector encoding for two example expressions.
+4 LLM-Informed Finite Expression Method
+In this section, we integrate the predictive operator set extracted from PDE data, namely the forcing $f$, boundary data $g$, and geometric/metrical descriptors of $\partial\Omega$, into an efficient symbolic-learning framework for solving PDEs: the finite-expression (FEX) method [32].
+The LLM-informed FEX pipeline for solving PDEs is illustrated in Fig.
+4
+Figure 4: LLM-informed FEX for interpretable PDE solutions.
+FEX poses the construction of fully interpretable PDE solutions as a reinforcement learning search over an operator library; a policy samples symbolic programs whose execution yields an explicit surrogate $u_{\theta}$, and the return is a physics-informed score that penalizes interior residuals and boundary mismatches, as defined in (2).
+Although FEX has demonstrated strong practical performance for high-dimensional PDEs, the convergence behavior of its policy-gradient optimization has remained unclear.
+We close this gap by establishing conditions under which policy-gradient iterates converge to stationary policies, thereby providing the novel convergence guarantees for symbolic PDE learning with policy gradients and placing the overall pipeline on firm theoretical footing among state-of-the-art alternatives.
+After feeding the PDE data to the fine-tuned LLM, we obtain a predicted operator set that, together with a computational expression tree $\mathcal{T}$ (see Fig.
+1), induces a hypothesis class $\mathbb{S}$.
+We seek
+$u^{*}\in\operatorname*{arg\,min}_{u\in\mathbb{S}}L(u).$
+Writing $\mathbb{S}=\{u_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{m}\}$, each candidate $u\in\mathbb{S}$ is represented by a (possibly mixed discrete/continuous) parameter $\theta\in\Theta$.
+We model search over $\Theta$ with a stochastic policy $\theta\sim p_{\phi}$, where $p_{\phi}$ is a distribution on $\Theta$ parameterized by $\phi\in\Phi$ and $\Phi\subseteq\mathbb{R}^{n}$ is nonempty, closed, and convex.
+Defining $R(\theta):=(1+L(u_{\theta}))^{-1}\in[0,1]$, we can recast model selection as
+(6) $\max_{\phi\in\mathbb{R}^{n}}\;F(\phi)\;:=\;\mathbb{E}_{\theta\sim p_{\phi}}\!\big[R(\theta)\big]\;-\;\delta_{\Phi}(\phi),$
+with $\delta_{\Phi}$ the indicator of $\Phi$ (0 on $\Phi$, $\infty$ otherwise).
+We let $J(\phi):=\mathbb{E}_{\theta\sim p_{\phi}}[R(\theta)]$ for notational simplicity.
+Assuming $p_{\phi}$ has $\phi$-independent support, is differentiable in $\phi$, $R$ is integrable, and $\mathbb{E}_{p_{\phi}}\!\big[\|R(\theta)\nabla_{\phi}\log p_{\phi}(\theta)\|\big]<\infty$ (so that $\nabla$ and $\mathbb{E}$ interchange), the gradient of $J$ is easily obtained as follows:
+$\nabla J(\phi)=\mathbb{E}_{\theta\sim p_{\phi}}\!\big[R(\theta)\,\nabla\log p_{\phi}(\theta)\big].$
+Problem (6) is generally nonconvex, so we can only target the first-order stationarity, without assuming restricted additional conditions.
+In particular, a point $\phi\in\mathbb{R}^{n}$ is stationary if
+$0\in-\nabla J(\phi)+N_{\Phi}(\phi),$
+where $N_{\Phi}(\phi):=\{\,g\in\mathbb{R}^{n}:\langle g,\phi^{\prime}-\phi\rangle\leq 0,\ \forall\,\phi^{\prime}\in\Phi\,\}$ is the normal cone of $\Phi$ at $\phi$ [36].
+We call a (possibly random) iterate $\hat{\phi}$ an $\epsilon$-stationary point in expectation if, for a sequence $\{\phi_{t}\}_{t=0}^{T}$ generated by a stochastic optimization method and a standard output rule, i.e., uniformly at random from $\{1,\dots,T\}$, it holds that
+$\mathbb{E}_{T}\left[\mathrm{dist}\left(0,-\nabla J(\hat{\phi})+N_{\Phi}(\hat{\phi})\right)^{2}\right]\;\leq\;\epsilon^{2},$
+where the expectation is over all algorithmic randomness (policy sampling, minibatching, and the random output selection).
+With the above preparation, we can now apply the stochastic projected policy gradient method (SPPGM) for finding the optimal policy. The template of the SPPGM is presented in Algorithm 1.
+Algorithm 1 Stochastic Projected Policy Gradient Method (SPPGM)
+1:Feasible set $\Phi\subseteq\mathbb{R}^{n}$, projection $\Pi_{\Phi}(\cdot)$, initial parameter $\phi_{0}\in\Phi$, step sizes $\{\eta_{t}\}_{t\geq 0}$, batch size $B$
+2:for $t=0,1,2,\ldots,T-1$ do
+3:  Sampling: Draw i.i.d. $\theta_{1},\ldots,\theta_{B}\sim p_{\phi_{t}}$
+4:  Rewards: Evaluate $r_{i}\leftarrow R(\theta_{i})$ for $i=1,\ldots,B$
+5:  Score gradients: Compute $s_{i}\leftarrow\nabla_{\phi}\log p_{\phi}(\theta_{i})\big|_{\phi=\phi_{t}}$
+6:  Gradient estimate: $g_{t}\;\leftarrow\;\frac{1}{B}\sum_{i=1}^{B}r_{i}s_{i}$
+7:  Stochastic Gradient step: $\tilde{\phi}_{t+1}\leftarrow\phi_{t}+\eta_{t}\,g_{t}$
+8:  Projection: $\phi_{t+1}\leftarrow\Pi_{\Phi}\big(\tilde{\phi}_{t+1}\big) \triangleright$ $\Pi_{\Phi}(x):=\arg\min_{y\in\Phi}\;\tfrac{1}{2}\|y-x\|^{2}$
+9:  Stopping: if a criterion holds, break
+10:end for
+11:Output: $\hat{\phi}_{T}$ is selected uniformly at random from the generated sequence $\{\phi_{t}\}_{t=1}^{T}$
+We end this section by presenting the following theorem, which shows that the SPPGM is able to return an $\epsilon$-stationary point in expectation with a constant learning rate by assume that $J(\phi)$ is $L$-smooth, i.e., for any $\phi^{\prime},\phi$, there exists a constant $L>0$ such that
+$\|\nabla J(\phi^{\prime})-\nabla J(\phi)\|\leq L\|\phi^{\prime}-\phi\|.$
+It is well-known (see, e.g., [1]) that for any $L$-smooth function $J(\phi)$, it holds that
+(7) $J(\phi^{\prime})\geq J(\phi)+\nabla J(\phi)^{T}(\phi^{\prime}-\phi)-\frac{L}{2}\|\phi^{\prime}-\phi\|^{2},\quad\forall\;\phi^{\prime},\phi\in\mathbb{R}^{n}.$
+Theorem 4.1.
+Suppose that the function $J(\cdot)$ is $L$-smooth with $L>0$, $\eta_{t}=\eta\in(0,\tfrac{1}{2L})$ for all $t\geq 0$, and there exists a constant $\sigma$ such that
+$\mathbb{E}_{\theta\sim p_{\phi}}\left[\|R(\theta)\nabla\log p_{\phi}(\theta)-\nabla J(\theta)\|^{2}\right]\leq\sigma^{2}$
+then the Algorithm 1 outputs a point $\hat{\phi}_{T}\in\Phi$ satisfying
+(8) $\;\mathbb{E}_{T}\left[\mathrm{dist}\left(0,-\nabla J(\hat{\phi}_{T})+N_{\Phi}(\hat{\phi}_{T})\right)^{2}\right]$
+$\leq$ $\;\left(2+\frac{2}{\eta L(1-2\eta L)}\right)\frac{\sigma^{2}}{B}+\frac{\Delta}{T}\left(\frac{2}{\eta}+\frac{4}{\eta(1-2\eta L)}\right),$
+where $\Delta:=F^{*}-F(\phi_{0})>0$ with $F^{*}$ being the global minimum of the problem (6). Moreover, for any given $\epsilon>0$, if one chooses
+$B:=\left\lceil\frac{\sigma^{2}}{\epsilon^{2}}\left(4+\frac{4}{\eta L(1-2\eta L}\right)\right\rceil,\quad T=\left\lceil\frac{\Delta}{\epsilon^{2}}\left(\frac{4}{\eta}+\frac{8}{\eta(1-2\eta L)}\right)\right\rceil,$
+then $\mathbb{E}_{T}\left[\mathrm{dist}\left(0,-\nabla J(\hat{\phi}_{T})+N_{\Phi}(\hat{\phi}_{T})\right)^{2}\right]\leq\epsilon^{2}$, and the sample complexity is $\mathcal{O}\big(\epsilon^{-4}\big)$.
+Proof 4.2.
+From the $L$-smoothness of $J(\cdot)$ and (7), we see that
+(9) $J(\phi_{t+1})\geq J(\phi_{t})+\nabla J(\phi_{t})^{T}(\phi_{t+1}-\phi_{t})-\frac{L}{2}\|\phi_{t+1}-\phi_{t}\|^{2},\quad\forall t\geq 0.$
+Since $\phi_{t+1}=\Pi_{\Phi}(\phi_{t}+\eta g_{t})=\arg\min_{\phi\in\Phi}\,\tfrac{1}{2}\|\phi-(\phi_{t}+\eta g_{t})\|^{2}$, by the first-order optimality condition of this projection problem, it is easy to verify that
+(10) $g_{t}^{T}(\phi_{t+1}-\phi_{t})+\frac{1}{2\eta}\|\phi_{t+1}-\phi_{t}\|^{2}\leq 0,$
+and that
+(11) $g^{t}-\frac{1}{\eta}(\phi_{t+1}-\phi_{t})\in N_{\Phi}(\phi_{t+1}).$
+From (9) and (10), we have
+$J(\phi_{t+1})+g_{t}^{T}(\phi_{t+1}-\phi_{t})-\frac{1}{2\eta}\|\phi_{t+1}-\phi_{t}\|^{2}\geq J(\phi_{t})+\nabla J(\phi_{t})^{T}(\phi_{t+1}-\phi_{t})-\frac{L}{2}\|\phi_{t+1}-\phi_{t}\|^{2}.$
+Rearranging terms and using the fact that both $\phi_{t+1},\phi_{t}\in\Phi$ (hence, $\delta_{\Phi}(\phi_{t+1})=\delta_{\Phi}(\phi_{t})=0$), we can rewrite the above inequality as
+(12) $\frac{1-\eta L}{2\eta}\|\phi_{t+1}-\phi_{t}\|^{2}\leq F(\phi_{t+1})-F(\phi_{t})+(g_{t}-\nabla J(\phi_{t}))^{T}(\phi_{t+1}-\phi_{t}).$
+On the one hand, by the Cauchy-Schwarz inequality, we see that
+$(g_{t}-\nabla J(\phi_{t}))^{T}(\phi_{t+1}-\phi_{t})\leq\frac{1}{2L}\|g^{t}-\nabla J(\phi_{t})\|^{2}+\frac{L}{2}\|\phi_{t+1}-\phi_{t}\|^{2},$
+which together with (12) implies that
+$\frac{1-2\eta L}{2\eta}\|\phi_{t+1}-\phi_{t}\|^{2}\leq F(\phi_{t+1})-F(\phi_{t})+\frac{1}{2L}\|g^{t}-\nabla J(\phi_{t})\|^{2}.$
+Summing the above inequality over $t=0,\dots,T-1$, we get
+(13) $\;\frac{1-2\eta L}{2\eta}\sum_{t=0}^{T-1}\|\phi_{t+1}-\phi_{t}\|^{2}$
+$\leq$ $\;F(\phi_{T})-F(\phi_{0})+\frac{1}{2L}\sum_{t=0}^{T-1}\|g^{t}-\nabla J(\phi_{t})\|^{2}$
+$\leq$ $\;\Delta+\frac{1}{2L}\sum_{t=0}^{T-1}\|g^{t}-\nabla J(\phi_{t})\|^{2}.$
+On the other hand, adding the term $(\nabla J(\phi_{t+1})-g_{t})^{T}(\phi_{t+1}-\phi_{t})$ and then multiplying $\tfrac{2}{\eta}$ on both sides of (12), we obtain that
+(14)  $2(\nabla J(\phi_{t+1})-g_{t})^{T}\left(\frac{1}{\eta}(\phi_{t+1}-\phi_{t})\right)+\frac{1-\eta L}{\eta^{2}}\|\phi_{t+1}-\phi_{t}\|^{2}$
+$\leq$ $\;\frac{2}{\eta}(F(\phi_{t+1})-F(\phi_{t}))+\frac{2}{\eta}(\nabla J(\phi_{t+1})-\nabla J(\phi_{t}))^{T}(\phi_{t+1}-\phi_{t}).$
+Notice that
+ $2(\nabla J(\phi_{t+1})-g_{t})^{T}\left(\frac{1}{\eta}(\phi_{t+1}-\phi_{t})\right)$
+$=$ $\;\left\lVert\nabla J(\phi_{t+1})-g_{t}+\frac{1}{\eta}(\phi_{t+1}-\phi_{t})\right\rVert^{2}-\|\nabla J(\phi_{t+1})-g_{t}\|^{2}-\frac{1}{\eta^{2}}\|\phi_{t+1}-\phi_{t}\|^{2},$
+we see that (14) can be written as
+$\;\left\lVert\nabla J(\phi_{t+1})-g_{t}+\frac{1}{\eta}(\phi_{t+1}-\phi_{t})\right\rVert^{2}$
+$\leq$ $\;\|\nabla J(\phi_{t+1})-g_{t}\|^{2}+\left(\frac{1}{\eta^{2}}-\frac{1-\eta L}{\eta^{2}}\right)\|\phi_{t+1}-\phi_{t}\|^{2}+\frac{2}{\eta}(F(\phi_{t+1})-F(\phi_{t}))$
+$\;+\frac{2}{\eta}(\nabla J(\phi_{t+1})-\nabla J(\phi_{t}))^{T}(\phi_{t+1}-\phi_{t})$
+$\leq$  $2\|\nabla J(\phi_{t})-g_{t}\|^{2}+2\|\nabla J(\phi_{t+1})-\nabla J(\phi_{t})\|^{2}\quad\text{[$\|a-b\|^{2}\leq 2\|a-c\|^{2}+2\|b-c\|^{2}$]}$
+$\;+\frac{L}{\eta}\|\phi_{t+1}-\phi_{t}\|^{2}+\frac{2}{\eta}(F(\phi_{t+1})-F(\phi_{t}))$
+$\;+\frac{2}{\eta}\|\nabla J(\phi_{t+1})-\nabla J(\phi_{t})\|\|\phi_{t+1}-\phi_{t}\|\quad\text{[Cauchy-Schwarz inequality]}$
+$\leq$  $2\|\nabla J(\phi_{t})-g_{t}\|^{2}+\left(2L^{2}+\frac{3L}{\eta}\right)\|\phi_{t+1}-\phi_{t}\|^{2}\quad\text{[$L$-smoothness]}$
+$\;+\frac{2}{\eta}(F(\phi_{t+1})-F(\phi_{t})).$
+Summing this inequality over $t=0,\dots,T-1$, we get
+$\;\sum_{t=0}^{T-1}\left\lVert\nabla J(\phi_{t+1})-g_{t}+\frac{1}{\eta}(\phi_{t+1}-\phi_{t})\right\rVert^{2}$
+$\leq$  $2\sum_{t=0}^{T-1}\|\nabla J(\phi_{t})-g_{t}\|^{2}+\left(2L^{2}+\frac{3L}{\eta}\right)\sum_{t=0}^{T-1}\|\phi_{t+1}-\phi_{t}\|^{2}+\frac{2}{\eta}(F(\phi_{T})-F(\phi_{0}))$
+$\leq$  $2\sum_{t=0}^{T-1}\|\nabla J(\phi_{t})-g_{t}\|^{2}+\frac{2}{\eta^{2}}\sum_{t=0}^{T-1}\|\phi_{t+1}-\phi_{t}\|^{2}+\frac{2\Delta}{\eta}\quad\text{[$L<\frac{1}{2\eta}$]}$
+[…]
+$=$  $2\sum_{t=0}^{T-1}\|\nabla J(\phi_{t})-g_{t}\|^{2}+\frac{2}{\eta L(1-2\eta L)}\sum_{t=0}^{T-1}\|g^{t}-\nabla J(\phi_{t})\|^{2}+\Delta\left(\frac{2}{\eta}+\frac{4}{\eta(1-2\eta L)}\right)$
+$=$  $2\left(1+\frac{2}{\eta L(1-2\eta L)}\right)\sum_{t=0}^{T-1}\|\nabla J(\phi_{t})-g_{t}\|^{2}+\Delta\left(\frac{2}{\eta}+\frac{4}{\eta(1-2\eta L)}\right),$
+which further implies that
+$\;\mathbb{E}_{T}\left[\mathrm{dist}\left(0,-\nabla J(\hat{\phi}_{T})+N_{\Phi}(\hat{\phi}_{T})\right)^{2}\right]$
+$=$ $\;\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}_{T}\left[\mathrm{dist}\left(0,-\nabla J(\phi_{t+1})+N_{\Phi}(\phi_{t+1})\right)^{2}\right]\qquad\text{[output rule]}$
+$\leq$ $\;\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}_{T}\left[\left\lVert\nabla J(\phi_{t+1})-g_{t}+\frac{1}{\eta}(\phi_{t+1}-\phi_{t})\right\rVert^{2}\right]\qquad\text{[projection optimality]}$
+$\leq$ $\;\frac{1}{T}\left(2+\frac{2}{\eta L(1-2\eta L)}\right)\sum_{t=0}^{T-1}\mathbb{E}_{T}\left[\|\nabla J(\phi_{t})-g_{t}\|^{2}\right]+\frac{\Delta}{T}\left(\frac{2}{\eta}+\frac{4}{\eta(1-2\eta L)}\right)$
+$\leq$ $\;\frac{1}{T}\left(2+\frac{2}{\eta L(1-2\eta L)}\right)\sum_{t=0}^{T-1}\frac{\sigma^{2}}{B}+\frac{\Delta}{T}\left(\frac{2}{\eta}+\frac{4}{\eta(1-2\eta L)}\right)\quad\text{[mini-batch sampling]}$
+$=$ $\;\left(2+\frac{2}{\eta L(1-2\eta L)}\right)\frac{\sigma^{2}}{B}+\frac{\Delta}{T}\left(\frac{2}{\eta}+\frac{4}{\eta(1-2\eta L)}\right).$
+This proves (8) and the remaining statements follow as a direct consequence of (8). Hence, the proof is completed.
 5 Experimental Results
@@
 The uninformed FEX employs a fixed operator set:
-[…]
+$\text{Binary set: }\mathbb{B}=\{+,-,*\},$
+$\text{Unary set: }\mathbb{U}=\{0,1,\text{Id},(\cdot)^{2},(\cdot)^{3},(\cdot)^{4},\exp,\sin,\cos\}.$
 In contrast, the LLM-informed FEX dynamically predicts the operator sets for each example, enabling a more adaptive and potentially efficient search process.
 5.2 Effectiveness of Fine-tuned LLMs
-In this subsection, we evaluate the performance of the fine-tuned BART, T5 and Llama3 models in predicting operator sets for the two PDE types.
+In this subsection, we evaluate the performance of the fine-tuned BART, T5 and Llama3 models in predicting operator sets.
 Specifically, we track the average number of mismatched operators on the test dataset for each epoch.
 The computational results are summarized in Figure 5.
-Interested readers are referred to the Table 3 in Appendix C.2 for the detailed computational results.
 Figure 5: Comparison between T5, BART, Llama3-3B and Llama3-8B in terms of average mismatch on test dataset.
+Interested readers are referred to the Table 3 in Appendix A for the detailed computational results.
 Based on the computational results, LLaMA-8B consistently achieves the lowest number of average mismatch, starting at 0.83 and decreasing to 0.38 within 12 epochs, demonstrating superior performance and training stability.
@@
 Since FEX is a randomized reinforcement learning approach, each instance was solved five times to obtain average computational results, including the number of iterations until convergence, computation time, and solution error.
-The detailed computational results can be found in Appendix C.3, and Table 1 and Figure4 summarize the statistics of these results.
-PDE Type Metric Uninformed FEX LLM-Informed FEX Speedup
-Conservation Avg. Iter 28.52 5.23 5.45$\times$
+The detailed computational results can be found in Appendix B, and Table 2 and Figure 6 summarize the statistics of these results.
+PDE Metric Uninformed LLM-Informed Speedup
+Conserv Avg. Iter 28.52 5.23 $5.45\times$
 Avg. Time (m) 23.67 3.95 $6.00\times$
@@
 Avg. Time (m) 62.07 14.09 $4.41\times$
-Table 1: Comparison on computational efficiency.
+Table 2: Comparison of computational efficiency between uninformed and LLM-informed FEX.
 Figure 6: Comparison on solution accuracy. The accuracy is in log-domain.
@@
 These findings demonstrate that integrating the fine-tuned LLM into symbolic machine learning based PDE solvers enables more efficiency and reliability via informed operator selection.
+Future research directions include developing a unified theoretical framework for a broader class of PDEs, conducting extensive numerical experiments on complex real-world problems to evaluate practical effectiveness, and benchmarking existing symbolic regression methods or designing novel frameworks tailored to PDE solution discovery.
 6 Conclusions
@@
 These findings highlight the promising role of LLMs in advancing symbolic reasoning for scientific problems, paving the way for further exploration of AI-driven methodologies in mathematical modeling and equation solving.
-Acknowledgments and Disclosure of Funding
-The authors were partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and the DARPA D24AP00325-00.
-Appendix A Binary Expression Trees
-The building blocks of the FEX methods in [30] are illustrated in the following Table 5.
-By increasing the depth of the rexpression tree, the number of functions that can be represented can be significantly increased.
-Also, the mechanism can automatically handle high-dimensional variables.
-Figure 5: Computational expression tree [30].
-Appendix B Proof of Theorem 1
-Let $W_{t},\;t\geq 0$ be a $d$-dimensional Brownian motion such that $W_{t}$ is $F_{t}$-measurable with probability measure $P$. We define the stochastic process
-$X_{t}=X_{0}+W_{t},\quad t\geq 0,$
-and denote by $P_{x}$ the probability measure conditioned on $X_{0}=x$ for any $x\in\mathbb{R}^{d}$, and the expectation with respect to $P_{x}$ is denoted as $\mathbb{E}_{x}[\cdot]$.
-Accordingly, we define the discrete process $\bar{X}_{k},\;k\geq 0$ as follows:
-$\bar{X}_{k}=\bar{X}_{k-1}+Y_{k}\mathcal{D}(\bar{X}_{k-1}),\quad k\geq 1,\quad X_{0}=x,$
-where $Y_{k},\;k\geq 1$ is indpendently, indentically and uniformly distribution on the unit sphere according to $X_{\tau(B(0,1))}$ with respect to a probability distribution $P_{0}$.
-Obviously, $\bar{X}_{k}$ depends on $x,Y_{1},\dots,Y_{k}$, hence to emphasize this dependency, we also dentoe
-$\bar{X}_{k}:=\bar{X}_{k}(x,Y_{1},\dots,Y_{k}),\quad k\geq 1.$
-For any open, nonempty set $\mathcal{X}\subseteq\mathbb{R}^{d}$, we define the first exit time of the process $X_{t}$ starting at $x\in\mathcal{X}$ from $\mathcal{X}$ by
-$\tau(\mathcal{X}):=\inf\left\{t\geq 0\;:\;X_{t}\notin\mathcal{X}\right\}.$
-Then, from [23], we see that for $x\in\Omega$, the solution $u(x)$ for the Poisson equation admits the following expression:
-[…]
-We first consider approximating the first term in (5) using the Monte-Carlo sampling technique. To this end, we define the function $g^{(M,\{K_{i}\})}:\mathbb{R}^{d}\to\mathbb{R}$ as
-$g^{(M,\{K_{i}\})}(x):=\frac{1}{M}\sum_{i=1}^{M}g\left(\bar{X}_{K_{i}}(x,Y_{i,1},\dots,Y_{i,K_{i}})\right),$
-where $M\geq 1$, $K_{i}\geq 1$ for $i=1,\dots,M$, and $Y_{i,k},\;k=1,\dots,K_{i}$ are unit vectors.
-Then, for any $\delta\in(0,1)$, there exist $M$ and $K_{i}$, $1\leq i\leq M$, all polynomial in $\delta^{-1}$ [15], such that
-[…]
-Similarly, we consider approximating the second term in (5) using the Monte-Carlo sampling technique.
-Let $Y_{i,k},\;k=1,\dots,K_{i}$ and $y_{i,j,k},\;k=1,\dots,K_{i},\;j=1,\dots,J$ be random unit vectors, for $i=1,\dots,M$, with $M,K_{i},J$ being positive integers, then for any $\delta>0$, one can show that the following function $f^{(M,\{K_{i}\},J)}:\mathbb{R}^{d}\to\mathbb{R}$ defined by
-$f^{(M,\{K_{i}\},J)}(x):=$ $\;\frac{1}{MJ}\sum_{i=1}^{M}\sum_{k=1}^{M_{i}}\sum_{j=1}^{J}\mathcal{D}\left(\bar{X}_{k}(x,Y_{i,1},\dots,Y_{i,k})\right)^{2}$
-×$\qquad\times f\left(\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1})+\mathcal{D}(\bar{X}_{k-1}(x,Y_{i,1},\dots,Y_{i,k-1}))y_{i,j,k}\right),$
-can approximate the second term (5) in sense that
-[…]
-if one chooses suitable $M,\;J$ and $K_{i},\;i=1,\dots,M$ all polynomial in $\delta^{-1}$ [15]. Therefore, the proof is completed.
-Appendix C Detailed Computational Results
-C.1 Illustrative Example for Binary Encoding
-Consider the following two functions:
-$h_{1}(x_{1},x_{2}):=(5x_{1})^{2}+\sin(3x_{2})\cdot x_{2},\quad h_{2}(x_{1}):=5\exp(2x_{1})\cdot(\cos(6x_{1}))^{3}.$
-The binary representation of them is provided in Table 2.
-Binary Vector
-Function Preprocessed Operator Set
-$x_{1}$
-$x_{2}$
-^2
-^3
-SIN
-COS
-EXP
-$h_{1}(x_{1},x_{2})$ [ $x_{1}$, $x_{2}$, ^2, +, *, SIN ] 1 1 1 0 1 1 1 0 0
-$h_{2}(x_{1})$ [ $x_{1}$, ^3, *, COS, EXP ] 1 0 0 1 0 1 0 1 1
-Table 2: Sample binary-vector encoding for two example expressions. The size of the dictionary is 9.
-The resulted mismatch can then be computed as $7$, indicating that the two functions are very different from each other.
-C.2 Detailed Computational Results for Average Mismatch
+Acknowledgments
+The authors were partially supported by the US National Science Foundation under awards DMS-2244988, IIS-25209787, GEO-5239902, and the DARPA D24AP00325-00.
+Appendix A Detailed Computational Results for Average Mismatch
+See Table 3.
 Epoch LLaMA-8B LLaMA-3B T5 BART
@@
 Table 3: Comparison of Average Mismatch Across Models During Fine-tuning on 198K Dataset
-C.3 Detailed computational results for FEX
-True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $2\bigl{(}-2x_{2}^{3}+2\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\exp\bigl{(}4x_{2}\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8x_{2}^{4}-2\bigl{(}2\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4x_{2}^{4}+4x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\exp(x_{1})+2\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\sin\bigl{(}2x_{3}^{3}-2\exp(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4x_{2}^{4}+4x_{3}^{4}+8\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\sin\bigl{(}2x_{3}^{3}+2\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\exp\bigl{(}2x_{2}^{2}-2\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\sin^{2}(x_{2})\cos^{2}(x_{1})$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\exp\bigl{(}2\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\sin^{4}(x_{1})\cos^{4}(x_{1})$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\exp(2x_{2})-2\exp\bigl{(}2x_{1}^{2}-2x_{2}^{2}\bigr{)}$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\exp\bigl{(}2x_{3}^{4}-2\exp(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2x_{2}^{4}+2x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2x_{2}^{4}-2x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\sin\bigl{(}2\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\exp\bigl{(}2\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\cos\bigl{(}4x_{2}^{4}\bigr{)}\cos\bigl{(}2x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\sin\bigl{(}4\exp(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2x_{3}^{4}-2\bigr{)}^{3}+2\cos\bigl{(}2x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{4}+2\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $22\sin\bigl{(}2x_{1}^{4}-2\sin(x_{3})\bigr{)}-$ 2 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $22\sin\bigl{(}2x_{3}+2\bigr{)}-$ 2 2.2 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\cos\bigl{(}2\sin(x_{2})+2\cos(x_{3})\bigr{)}$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\cos\bigl{(}2x_{1}^{3}-2x_{3}^{2}\bigr{)}$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{2}+2\sin(x_{3})\bigr{)}+$ 2 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $232x_{3}^{6}+$ 2 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$ $2\bigl{(}-2x_{2}^{3}+2\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\exp\bigl{(}4x_{2}\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8x_{2}^{4}-2\bigl{(}2\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4x_{2}^{4}+4x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\exp(x_{1})+2\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\sin\bigl{(}2x_{3}^{3}-2\exp(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4x_{2}^{4}+4x_{3}^{4}+8\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\sin\bigl{(}2x_{3}^{3}+2\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\exp\bigl{(}2x_{2}^{2}-2\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\sin^{2}(x_{2})\cos^{2}(x_{1})$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\exp\bigl{(}2\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\sin^{4}(x_{1})\cos^{4}(x_{1})$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\exp(2x_{2})-2\exp\bigl{(}2x_{1}^{2}-2x_{2}^{2}\bigr{)}$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\exp\bigl{(}2x_{3}^{4}-2\exp(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2x_{2}^{4}+2x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2x_{2}^{4}-2x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\sin\bigl{(}2\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\exp\bigl{(}2\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\cos\bigl{(}4x_{2}^{4}\bigr{)}\cos\bigl{(}2x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\sin\bigl{(}4\exp(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2x_{3}^{4}-2\bigr{)}^{3}+2\cos\bigl{(}2x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $22\exp\bigl{(}2x_{1}^{4}+2\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$
+Appendix B Detailed computational results for FEX
+True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $2\bigl(-2x_{2}^{3}+2\sin(x_{3})\bigr)^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\exp\bigl(4x_{2}\cos(x_{2})\bigr)$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8x_{2}^{4}-2\bigl(2\cos(x_{2})+2\bigr)^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $4x_{2}^{4}+4x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl(-2\exp(x_{1})+2\cos(x_{1})\bigr)^{2}$ 38.0 7.8 82.448 15.878 $88\times 10^{-3} 2.2$$21\times 10^{-3}$ $2\sin\bigl(2x_{3}^{3}-2\exp(x_{3})\bigr)$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $4x_{2}^{4}+4x_{3}^{4}+8\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\sin\bigl(2x_{3}^{3}+2\cos(x_{3})\bigr)$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\exp\bigl(2x_{2}^{2}-2\sin(x_{2})\bigr)$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\sin^{2}(x_{2})\cos^{2}(x_{1})$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\exp\bigl(2\sin(x_{2})+2\bigr)$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\sin^{4}(x_{1})\cos^{4}(x_{1})$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\exp(2x_{2})-2\exp\bigl(2x_{1}^{2}-2x_{2}^{2}\bigr)$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\exp\bigl(2x_{3}^{4}-2\exp(x_{3})\bigr)$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl(2x_{2}^{4}+2x_{2}^{3}\bigr)^{2}$ 17.6 6.0 28.374 11.458 $47\times 10^{-2} 1.1$$10\times 10^{-2}$ $2\bigl(2x_{2}^{4}-2x_{2}^{2}\bigr)^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $2\bigl(2x_{1}^{2}-2\bigr)^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $2\sin\bigl(2\sin(x_{2})-2\bigr)$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\exp\bigl(2\cos(x_{2})+2\bigr)$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\cos\bigl(4x_{2}^{4}\bigr)\cos\bigl(2x_{1}^{4}-2\bigr)$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\sin\bigl(4\exp(x_{1})\bigr)$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl(2x_{3}^{4}-2\bigr)^{3}+2\cos\bigl(2x_{2}-2\bigr)$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $2\exp\bigl(2x_{1}^{4}+2\sin(x_{3})\bigr)+2$ 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $2\sin\bigl(2x_{1}^{4}-2\sin(x_{3})\bigr)-2$ 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $\sin\bigl(2x_{3}+2\bigr)-2 2.2$ 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\cos\bigl(2\sin(x_{2})+2\cos(x_{3})\bigr)$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\cos\bigl(2x_{1}^{3}-2x_{3}^{2}\bigr)$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ $2\exp\bigl(2x_{1}^{2}+2\sin(x_{3})\bigr)+2$ 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $32x_{3}^{6}+2$ 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$ $2\bigl(-2x_{2}^{3}+2\sin(x_{3})\bigr)^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\exp\bigl(4x_{2}\cos(x_{2})\bigr)$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8x_{2}^{4}-2\bigl(2\cos(x_{2})+2\bigr)^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $4x_{2}^{4}+4x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl(-2\exp(x_{1})+2\cos(x_{1})\bigr)^{2}$ 38.0 7.8 82.448 15.878 $88\times 10^{-3} 2.2$$21\times 10^{-3}$ $2\sin\bigl(2x_{3}^{3}-2\exp(x_{3})\bigr)$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $4x_{2}^{4}+4x_{3}^{4}+8\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\sin\bigl(2x_{3}^{3}+2\cos(x_{3})\bigr)$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\exp\bigl(2x_{2}^{2}-2\sin(x_{2})\bigr)$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\sin^{2}(x_{2})\cos^{2}(x_{1})$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\exp\bigl(2\sin(x_{2})+2\bigr)$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\sin^{4}(x_{1})\cos^{4}(x_{1})$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\exp(2x_{2})-2\exp\bigl(2x_{1}^{2}-2x_{2}^{2}\bigr)$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\exp\bigl(2x_{3}^{4}-2\exp(x_{3})\bigr)$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl(2x_{2}^{4}+2x_{2}^{3}\bigr)^{2}$ 17.6 6.0 28.374 11.458 $47\times 10^{-2} 1.1$$10\times 10^{-2}$ $2\bigl(2x_{2}^{4}-2x_{2}^{2}\bigr)^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $2\bigl(2x_{1}^{2}-2\bigr)^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $2\sin\bigl(2\sin(x_{2})-2\bigr)$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\exp\bigl(2\cos(x_{2})+2\bigr)$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\cos\bigl(4x_{2}^{4}\bigr)\cos\bigl(2x_{1}^{4}-2\bigr)$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\sin\bigl(4\exp(x_{1})\bigr)$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl(2x_{3}^{4}-2\bigr)^{3}+2\cos\bigl(2x_{2}-2\bigr)$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $2\exp\bigl(2x_{1}^{4}+2\sin(x_{3})\bigr)+2$ 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$
 Table 4: Results of LLM-informed and original FEX runs on 50 randomly generated Linear Conservation Law functions.
-True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $8\,x_{1}\,\mathrm{exp}(x_{2})$ 42.4 1.2 74.44 0.5900 $3.75\times 10^{-4}$ $4.91\times 10^{-4}$ $2128\,x_{1}^{12}+$ 2 40.2 1.2 70.704 0.516 $3.97\times 10^{-2}$ $2.14\times 10^{-2}$ $-32\,x_{2}^{10}$ 35.2 1.2 71.586 0.536 $7.44\times 10^{-3}$ $6.71\times 10^{-3}$ $-8\,x_{1}^{4}\,x_{2}$ 25.0 1.2 53.342 0.510 $9.18\times 10^{-4}$ $3.86\times 10^{-4}$ $2\,\cos\bigl{(}4\,x_{2}^{6}\bigr{)}$ 44.2 3.0 85.792 4.326 $1.68\times 10^{-2}$ $1.23\times 10^{-2}$ $-8\,x_{2}\,\cos(x_{2})+2$ 32.4 2.4 78.076 4.548 $3.78\times 10^{-4}$ $3.06\times 10^{-4}$ $28\,x_{1}^{2}\,\cos(x_{2})+$ 2 42.8 3.2 85.55 6.154 $2.05\times 10^{-4}$ $1.85\times 10^{-4}$ $-4\,\mathrm{exp}(x_{3})+4\,\sin(x_{1})+2$ 20.2 1.6 43.712 2.150 $1.14\times 10^{-4}$ $2.04\times 10^{-4}$ 2($22\bigl{(}2\,x_{2}^{2}-2\,x_{3}^{4}\bigr{)}^{2}+$ 2 31.8 3.0 69.054 3.400 $6.55\times 10^{-3}$ $2.66\times 10^{-3}$ $-32\,\sin(x_{2})^{4}$ 15.6 1.6 36.210 1.704 $4.34\times 10^{-3}$ $4.66\times 10^{-3}$ $-4\,x_{2}^{3}-4\,x_{3}^{4}+16$ 20.6 2.4 45.110 2.716 $3.88\times 10^{-4}$ $4.68\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}-2\,\mathrm{exp}(x_{3})+2\,\cos(x_{3})\bigr{)}+$ 2 22.0 2.8 52.848 4.994 $3.83\times 10^{-4}$ $4.28\times 10^{-4}$ $28\,x_{1}^{2}\,x_{3}+$ 2 11.4 1.6 26.022 1.742 $3.58\times 10^{-4}$ $1.82\times 10^{-4}$ $2\,\sin\bigl{(}2\,\sin(x_{1})-2\,\cos(x_{3})\bigr{)}$ 19.2 3.0 47.168 4.768 $7.78\times 10^{-5}$ $3.13\times 10^{-5}$ $4\bigl{(}2\,\cos(x_{2})-2\,\cos(x_{3})\bigr{)}^{4}$ 10.2 1.6 21.852 1.476 $3.80\times 10^{-3}$ $3.12\times 10^{-3}$ 2($22\bigl{(}2\,x_{1}+2\,\mathrm{exp}(x_{1})\bigr{)}^{2}-$ 2 26.4 4.2 44.664 6.108 $1.07\times 10^{-3}$ $2.94\times 10^{-3}$ $22\,\sin\bigl{(}2\,x_{2}^{3}+2\,x_{3}^{4}\bigr{)}+$ 2 29.0 4.8 73.818 12.376 $6.12\times 10^{-4}$ $3.11\times 10^{-4}$ 2($22\bigl{(}-2\,x_{1}+2\,\mathrm{exp}(x_{3})\bigr{)}^{2}-$ 2 20.4 3.4 35.796 4.762 $1.96\times 10^{-3}$ $2.52\times 10^{-3}$ $22\,\cos\bigl{(}2\,\mathrm{exp}(x_{1})+2\,\sin(x_{2})\bigr{)}+$ 2 69.6 11.6 157.764 27.654 $3.92\times 10^{-5}$ $1.17\times 10^{-4}$ $2\,\sin\bigl{(}2\,\sin(x_{2})-2\,\cos(x_{2})\bigr{)}$ 18.2 3.2 37.792 5.254 $1.01\times 10^{-5}$ $3.35\times 10^{-5}$ $2\bigl{(}-2\,x_{2}^{3}+2\,\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\,\mathrm{exp}\bigl{(}4\,x_{2}\,\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8\,x_{2}^{4}-2\bigl{(}2\,\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4\,x_{2}^{4}+4\,x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\,\mathrm{exp}(x_{1})+2\,\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\,\sin\bigl{(}2\,x_{3}^{3}-2\,\mathrm{exp}(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4\,x_{2}^{4}+4\,x_{3}^{4}+8\,\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\,\sin\bigl{(}2\,x_{3}^{3}+2\,\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\,\mathrm{exp}\bigl{(}2\,x_{2}^{2}-2\,\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\,\sin(x_{2})^{2}\,\cos(x_{1})^{2}$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\,\mathrm{exp}\bigl{(}2\,\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\,\sin(x_{1})^{4}\,\cos(x_{1})^{4}$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\,\mathrm{exp}(2\,x_{2})-2\,\mathrm{exp}(2\,x_{1}^{2}-2\,x_{2}^{2})$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\,\mathrm{exp}\bigl{(}2\,x_{3}^{4}-2\,\mathrm{exp}(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2\,x_{2}^{4}+2\,x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2\,x_{2}^{4}-2\,x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2\,x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\,\sin\bigl{(}2\,\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\,\mathrm{exp}\bigl{(}2\,\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\,\cos\bigl{(}4\,x_{2}^{4}\bigr{)}\,\cos\bigl{(}2\,x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\,\sin\bigl{(}4\,\mathrm{exp}(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2\,x_{3}^{4}-2\bigr{)}^{3}+2\,\cos\bigl{(}2\,x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\,\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}2\,x_{1}^{4}+2\,\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $22\,\sin\bigl{(}2\,x_{1}^{4}-2\,\sin(x_{3})\bigr{)}-$ 2 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $22\,\sin\bigl{(}2\,x_{3}+2\bigr{)}-$ 2 2.2 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\,\cos\bigl{(}2\,\sin(x_{2})+2\,\cos(x_{3})\bigr{)}$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\,\cos\bigl{(}2\,x_{1}^{3}-2\,x_{3}^{2}\bigr{)}$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}2\,x_{1}^{2}+2\,\sin(x_{3})\bigr{)}+$ 2 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $232\,x_{3}^{6}+$ 2 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$
+True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $8\,x_{1}\,\mathrm{exp}(x_{2})$ 42.4 1.2 74.44 0.5900 $3.75\times 10^{-4}$ $4.91\times 10^{-4}$ $128\,x_{1}^{12}+2$ 40.2 1.2 70.704 0.516 $3.97\times 10^{-2}$ $2.14\times 10^{-2}$ $32\,x_{2}^{10}$ 35.2 1.2 71.586 0.536 $7.44\times 10^{-3}$ $6.71\times 10^{-3}$ $8\,x_{1}^{4}\,x_{2}$ 25.0 1.2 53.342 0.510 $9.18\times 10^{-4}$ $3.86\times 10^{-4}$ $2\,\cos\bigl(4\,x_{2}^{6}\bigr)$ 44.2 3.0 85.792 4.326 $1.68\times 10^{-2}$ $1.23\times 10^{-2}$ $8\,x_{2}\,\cos(x_{2})+2$ 32.4 2.4 78.076 4.548 $3.78\times 10^{-4}$ $3.06\times 10^{-4}$ $8\,x_{1}^{2}\,\cos(x_{2})+2$ 42.8 3.2 85.55 6.154 $2.05\times 10^{-4}$ $1.85\times 10^{-4}$ $4\,\mathrm{exp}(x_{3})+4\,\sin(x_{1})+2$ 20.2 1.6 43.712 2.150 $1.14\times 10^{-4}$ $2.04\times 10^{-4}$ $2\bigl(2\,x_{2}^{2}-2\,x_{3}^{4}\bigr)^{2}+2$ 31.8 3.0 69.054 3.400 $6.55\times 10^{-3}$ $2.66\times 10^{-3}$ $32\,\sin(x_{2})^{4}$ 15.6 1.6 36.210 1.704 $4.34\times 10^{-3}$ $4.66\times 10^{-3}$ $4\,x_{2}^{3}-4\,x_{3}^{4}+16$ 20.6 2.4 45.110 2.716 $3.88\times 10^{-4}$ $4.68\times 10^{-4}$ $\mathrm{exp}\bigl(-2\,\mathrm{exp}(x_{3})+2\,\cos(x_{3})\bigr)+2 22$.0 2.8 52.848 4.994 $3.83\times 10^{-4}$ $4.28\times 10^{-4}$ $8\,x_{1}^{2}\,x_{3}+2$ 11.4 1.6 26.022 1.742 $3.58\times 10^{-4}$ $1.82\times 10^{-4}$ $2\,\sin\bigl(2\,\sin(x_{1})-2\,\cos(x_{3})\bigr)$ 19.2 3.0 47.168 4.768 $7.78\times 10^{-5}$ $3.13\times 10^{-5}$ $4\bigl(2\,\cos(x_{2})-2\,\cos(x_{3})\bigr)^{4}$ 10.2 1.6 21.852 1.476 $3.80\times 10^{-3}$ $3.12\times 10^{-3}$ $2\bigl(2\,x_{1}+2\,\mathrm{exp}(x_{1})\bigr)^{2}-2$ 26.4 4.2 44.664 6.108 $1.07\times 10^{-3}$ $2.94\times 10^{-3}$ $2\,\sin\bigl(2\,x_{2}^{3}+2\,x_{3}^{4}\bigr)+2$ 29.0 4.8 73.818 12.376 $6.12\times 10^{-4}$ $3.11\times 10^{-4}$ $2\bigl(-2\,x_{1}+2\,\mathrm{exp}(x_{3})\bigr)^{2}-2$ 20.4 3.4 35.796 4.762 $1.96\times 10^{-3}$ $2.52\times 10^{-3}$ $2\,\cos\bigl(2\,\mathrm{exp}(x_{1})+2\,\sin(x_{2})\bigr)+2$ 69.6 11.6 157.764 27.654 $3.92\times 10^{-5}$ $1.17\times 10^{-4}$ $2\,\sin\bigl(2\,\sin(x_{2})-2\,\cos(x_{2})\bigr)$ 18.2 3.2 37.792 5.254 $1.01\times 10^{-5}$ $3.35\times 10^{-5}$ $2\bigl(-2\,x_{2}^{3}+2\,\sin(x_{3})\bigr)^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\,\mathrm{exp}\bigl(4\,x_{2}\,\cos(x_{2})\bigr)$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8\,x_{2}^{4}-2\bigl(2\,\cos(x_{2})+2\bigr)^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $4\,x_{2}^{4}+4\,x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl(-2\,\mathrm{exp}(x_{1})+2\,\cos(x_{1})\bigr)^{2}$ 38.0 7.8 82.448 15.878 $88\times 10^{-3} 2.2$$21\times 10^{-3}$ $2\,\sin\bigl(2\,x_{3}^{3}-2\,\mathrm{exp}(x_{3})\bigr)$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $4\,x_{2}^{4}+4\,x_{3}^{4}+8\,\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\,\sin\bigl(2\,x_{3}^{3}+2\,\cos(x_{3})\bigr)$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\,\mathrm{exp}\bigl(2\,x_{2}^{2}-2\,\sin(x_{2})\bigr)$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\,\sin(x_{2})^{2}\,\cos(x_{1})^{2}$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\,\mathrm{exp}\bigl(2\,\sin(x_{2})+2\bigr)$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\,\sin(x_{1})^{4}\,\cos(x_{1})^{4}$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\,\mathrm{exp}(2\,x_{2})-2\,\mathrm{exp}(2\,x_{1}^{2}-2\,x_{2}^{2})$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\,\mathrm{exp}\bigl(2\,x_{3}^{4}-2\,\mathrm{exp}(x_{3})\bigr)$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl(2\,x_{2}^{4}+2\,x_{2}^{3}\bigr)^{2}$ 17.6 6.0 28.374 11.458 $47\times 10^{-2} 1.1$$10\times 10^{-2}$ $2\bigl(2\,x_{2}^{4}-2\,x_{2}^{2}\bigr)^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $2\bigl(2\,x_{1}^{2}-2\bigr)^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $2\,\sin\bigl(2\,\sin(x_{2})-2\bigr)$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\,\mathrm{exp}\bigl(2\,\cos(x_{2})+2\bigr)$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\,\cos\bigl(4\,x_{2}^{4}\bigr)\,\cos\bigl(2\,x_{1}^{4}-2\bigr)$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\,\sin\bigl(4\,\mathrm{exp}(x_{1})\bigr)$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl(2\,x_{3}^{4}-2\bigr)^{3}+2\,\cos\bigl(2\,x_{2}-2\bigr)$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\,\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $2\,\mathrm{exp}\bigl(2\,x_{1}^{4}+2\,\sin(x_{3})\bigr)+2$ 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $2\,\sin\bigl(2\,x_{1}^{4}-2\,\sin(x_{3})\bigr)-2$ 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $\sin\bigl(2\,x_{3}+2\bigr)-2 2.2$ 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\,\cos\bigl(2\,\sin(x_{2})+2\,\cos(x_{3})\bigr)$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\,\cos\bigl(2\,x_{1}^{3}-2\,x_{3}^{2}\bigr)$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ $2\,\mathrm{exp}\bigl(2\,x_{1}^{2}+2\,\sin(x_{3})\bigr)+2$ 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $32\,x_{3}^{6}+2$ 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$
 Table 5: Results of LLM-informed and original FEX runs on 50 randomly generated Poisson functions.